### v1

# Adaptive and robust watermark against model extraction attack

Abstract
Large language models have boosted Large Models as a Service (LMaaS) into a thriving business sector. But even model owners offering only API access while keeping model parameters and internal workings private, their Intellectual Property (IP) are still at risk of theft through model extraction attacks. To safeguard the IP of these models and mitigate unfair competition in the language model market, watermarking technology serves as an efficient post-hoc solution for identifying IP infringements. However, existing IP protection watermarking methods either explicitly alter the original output of the language model or implant watermark signals in the model logits. These methods forcefully distort the original distribution of the language model and impact the sampling process, leading to a decline in the quality of the generated text. The existing method also fails to achieve end-to-end adaptive watermark embedding and lack robustness verification in complex scenarios where watermark detection is subject to interference. To overcome these challenges, we propose PromptShield, a plug-and-play IP protection watermarking method to resist model extraction attacks without training additional modules. Leveraging the self-reminding properties inherent in large language models, we encapsulate the user’s query with a watermark self-generated instruction, nudging the LLMs to automatically generate watermark words in its output without compromising generation quality. Our method does not require access to the model’s internal logits and minimizes alterations to the model’s distribution using prompt-guided cues. Comprehensive experimental results consistently demonstrate the effectiveness, harmlessness, and robustness of our watermark. Moreover, Our watermark detection method remains robust and high detection sensitivity even when subjected to interference.
Index Terms:
Large Language Models, Model Extraction Attack, Watermarking
I Introduction
In the current era of Language Model as a Service (LMaaS), safeguarding the intellectual property (IP) of large language models (LLMs) becomes a critical issue. The training of LLMs demands significant intellectual efforts, involving massive amounts of high-quality data, extensive computational resources, and human-elaborated training design. These investments imbue LLMs with substantial knowledge value, necessitating their inclusion under the IP rights of the legal owner.
Figure 1: Language models offering services are at risk of model extraction attacks, which unfold in three steps, depicted in the diagram with blue boxes. Attackers first collect data output by the victim model through queries, then use these data to train their own imitation models. Eventually, the imitation model achieves performance comparable to the victim model, significantly endangering the IP of the victim model. IP protection watermarking involves embedding a special watermark signal in the output of the victim model. When watermarked data are used to train an imitation model, the watermark signal can still be detected in the model’s output. The stages of watermark generation and extraction are highlighted in yellow boxes in the diagram.
Despite the rapid growth of the LMaaS market, the development of comprehensive legal frameworks and regulations for IP protection lags behind [1, 2, 3]. The service process of LLMs is vulnerable to intellectual property infringement through model extraction attacks, also referred to as model stealing or imitation attacks. Model extraction attacks are akin to knowledge distillation [4], illustrated by the Alpaca case [5]. These attacks create a functionally comparable model in specific domains by distilling the victim model’s knowledge based on the queried victim model’s outputs [6, 7]. Such practices, especially when victim LLMs’ outputs are used for commercial model training, constitute a clear IP infringement. OpenAI’s policy [^1] Terms of use: https://openai.com/policies/terms-of-use, for instance, prohibits using ChatGPT’s outputs for competitive model training, underscoring the legal complexities and the need for stringent IP protection in AI development. Some top AI companies even have been caught stealing their competitors’ model outputs for training [8].
To avoid language model IP copyrights being offended, several real-time proactive methods for detecting attackers [9, 10, 11] and post-attack verification methods [12, 13, 14, 15, 16] for ownership authentication have been proposed. Watermarking, an information-hiding technique that embeds secret messages into carrier data, has been widely used for ownership protection [17, 18, 19, 20]. The purpose of watermarking is not to prevent IP stealing, as we cannot stop the attacker from imitating the actions of a normal user. Watermarks serve as a passive verification method, enabling the post-theft verification of IP ownership and complementing proactive defense strategies.
Existing IP protection watermarking methods typically involve embedding special signals into the output of the protected model as evidence to identify the suspect model’s IP ownership by designing heuristic watermark rules [15, 14, 16]. For example, He [15] employs lexical modification to embed a watermark in the output text of a targeted model. This watermarked data can then be assimilated through imitation model learning. Such token replacement visibly alters the original language model’s prediction results at the token level. Zhao [16] implements a unique cosine signal to manipulate the output logits of the model. The logits of every token at each timestep are adjusted to conform to a specific grouped cosine pattern, which the imitation model can then learn. This method also modifies the original distribution of the language model at the token probabilities level. Most existing IP protection watermarks are designed by modifying the distribution of language models through heuristic watermarking rules to obtain watermarking space. However, this inevitably distorts the original distribution of the language model and impacts the sampling process, resulting in the final sample deviating from optimal sampling and subsequently affecting the model’s performance in its downstream tasks. Also, these heuristic watermarking rules also lack sufficient flexibility and adaptability in real-world scenarios, struggling with unpredictable user queries that result in issues either outside the range of the watermark word list or with entropy too low to be effectively grouped.
Moreover, there is a lack of robust watermark detection in complex scenarios. Once attackers have obtained responses from the victim model in real-world scenarios, they can launch unpredictable and diverse attacks on the watermarked data. Attackers can also edit or replace watermarked text, which may nullify watermarks that rely on fixed synonym replacement rules. Discussions in content watermarking [21, 22] often highlight edit attacks at both the text token-level and sentence-level. Watermark robustness in scenarios of anti-model extraction attacks should also be resilient against potential dilution processes by attackers. Attackers might train imitation models using only a subset of query data or incorporate unwatermarked data. However, most existing IP protection watermarking methods have not considered the robustness of the designed watermarks, or have only accounted for robustness against specific attacks. This oversight is evident both in the lack of robustness in the watermark design mechanism and the limited empirical robustness demonstrated in experiments.
To address these drawbacks, we propose PromptShield, a simple but effective watermarking method for language model IP protection that avoids forcefully distorting the language model distribution and has minimal impact on the original task quality. We encapsulate the user’s query into the language model’s self-watermarking system prompt, which automatically cues the model to embed watermarks. Due to their stable instruction understanding abilities, LLMs can seamlessly incorporate watermarks while formulating responses. We initially explore the capability of large language models to self-watermark, and to our knowledge, we are the first to empower language models to autonomously generate watermarks. Our findings confirm the viability of this auto-generated watermarking method, underscored by the language models’ capacity to learn and incorporate these watermarks effectively.
Our contributions include:
  • We introduce a new plug-and-play adaptive watermarking method to combat model extraction attacks without additional training modules. Our method embeds watermarks automatically and directly during the inference stage without distorting language model distribution or disturbing the decoding, preserving the original performance of language models on downstream tasks.
  • We propose a robust watermark detection algorithm capable of effectively identifying watermarks even in realistic scenarios where the watermarks are subjected to interference. Remarkably, even if only 10 percent of watermarked data is utilized in training attackers’ models, our detector can confirm the unauthorized use of the protected model’s output.
  • Through extensive experimentation—including scenarios with different watermark detection strategies, different epochs, and robustness tests, we have validated the effectiveness, learnability, harmlessness, and robustness of our watermarks.
II Related Works
II-A Model Extraction Attack
Model extraction attacks [23, 24], especially targeting generative models, pose a serious security concern. The attacker aims to create a near-identical replica of the target model by using its outputs. Also known as imitation attacks or model stealing, attackers construct query-response pairs by accessing the victim model’s query API. Subsequently, these pairs are used to train a new model, allowing the imitation to achieve comparable performance to the victim model without needing knowledge of the internal details of the victim model, which renders many backdoor-based watermarking methods ineffective.
II-A1 Types of APIs
Model extraction attacks on victim language models vary based on the API service type. These attacks target label predictions, output embeddings, or output tokens.
Task-specific label-targeted methods are prevalent [25, 26, 27] at early stage. Attackers exploit the soft labels or hard labels returned by the API to train imitation models. For instance, Krishna [26] developed a model extraction method targeting BERT-APIs for NLP tasks such as sentiment classification and boolean question answering, utilizing the result probabilities as soft labels to train the imitation model. Sanyal [28] proposed a GAN-based framework that trains the imitation model and the generator in tandem to steal the victim model in a hard label setting.
During the era of Embedding as a Service (EaaS), attacks targeting embeddings became prevalent, facilitating a range of natural language processing tasks [24, 29, 30]. Embeddings can provide richer information than labels to assist in training imitation models. Liu [24] initiated the first attack strategy to exfiltrate pre-trained encoders by exploiting the embeddings returned from the victim model. Similarly, Shetty [29] introduced the CSE (Clustering, Selection, Elimination) framework to extract language models from cloud EaaS APIs.
However, the focus has largely shifted towards APIs that directly generate tokens outputs in the era of LLMs [31, 32]. Compared to embeddings, output tokens carry linguistic patterns and contextual dependencies. The imitation model can more easily learn the complex linguistic logic from these information. Wallace [31] launched an attack on a black-box translation system by inputting monolingual sentences and training models to mimic their outputs. Xu [32] demonstrated that black-box APIs, including those that only return tokens, are susceptible to model extraction. The attackers could even outperform the original victim models through unsupervised domain adaptation and multi-victim ensembles. Our research primarily focus on this output tokens-targeted category.
II-B Defense against Model Extraction Attack
To counter model extraction attacks, many efforts have been made to safeguard against model extraction attacks from different perspectives. We categorize defense strategies based on their timing of action into real-time defenses during querying and post-attack IP verification.
II-B1 Real-time defenses
Real-time defenses actively identify behaviors indicative of model theft. This is achieved by analyzing input queries [9, 11] or introducing perturbations to the model’s response [33, 34, 35], thereby complicating the theft process for attackers. PRADA [9] detects abnormal query patterns through statistical analysis of the distance distributions among queries, signaling potential model theft. Additionally, SEAT [11] utilizes a similarity encoder trained with adversarial techniques to spot pairs of similar queries, commonly employed by attackers as they progressively expand a set of seed samples to train the imitation model. Furthermore, defenders can even disseminate misinformation upon detecting out-of-distribution queries [10].
II-B2 post-attack defenses
In contrast, post-attack defenses concentrate on integrating identification markers, such as digital watermarks or fingerprints, into either the victim model or its output. Although post-attack defenses do not prevent the model from being stolen, it plays a pivotal role in asserting IP rights after theft [12, 13]. Watermarking offers a cost-effective and highly efficient approach to counter model extraction attacks. It also can be integrated with real-time defense mechanisms to enhance the overall defense strategy without requiring substantial computational resources. As shown in Fig 12 in appendix VI-A, we categorize watermarks into four types based on their protection targets: model parameter watermarks, content watermarks, dataset watermarks, and IP protection watermarks. We emphasize that the focus of our work is the IP protection for LLMs against model extraction attack, diverging from parameter watermarking [36, 37, 38, 39] to prevent the theft of the model parameters and content watermarking [21, 22, 40, 41] proving the ownership of model outputs. Specific introductions about the model parameter watermarks, content watermarks and dataset watermarks types, as well as the relationships between these watermarks, are presented in the appendix VI-A.
II-C Language model watermarking method against model extraction attack
The fundamental concept behind language model watermarking against model extraction attacks is to introduce special learnable signals into the outputs of the model to be protected. When attackers use these watermarked model outputs to train their imitation models, these special signals can be detected in the output of the imitation models, allowing for effective IP protection. Based on the watermarking method’s authority to access and manipulate internal computational parameters (such as logits), language model IP protection watermarking methods can be categorized into two distinct classes: black-box watermarking methods and white-box watermarking methods. The detailed workflow of these two types of methods is demonstrated in the appendix VI-B.
White-box methods require knowledge of the internal details of the model to be protected and the ability to manipulate the text generation process. The representative white-box method is GINSEW [16], which groups and manipulates the conditional token probabilities of language model. It embeds specific frequency sine perturbations as watermark signals in the grouped probabilities, enabling the imitation models to learn this signal pattern during training. Watermark detection involves querying the imitation model for token probabilities, conducting matching filtering to detect grouped conditional probability outputs, and checking for the presence of embedded specific frequency signals. Similarly, the EmbMarker [42] uses a set of medium-frequency trigger words as watermark signals. EmbMarker selects a target embedding as the watermark and inserts it into the embeddings of texts containing trigger words as a backdoor. However, it may not be suitable for the current LLM market compared to the previous era of EaaS.
Black-box methods, on the other hand, do not require knowledge of the internal details of the model to be protected and can provide more universal protection for language models. A classic work in the black-box domain is the lexical watermarking proposed by He [15] for protecting intellectual property. This method involves synonym replacement and English-American grammar substitution in the output of the language model to be protected. Specifically, a set of words $W$ is selected, and for each word $w$ in $W$, a semantically equivalent substitution is identified and used to replace $w$ in the output of the victim model. The objective is to alter the word distribution such that the imitation model learns this skewed pattern. Attackers, when training their own imitation models using the replaced data, tend to output these substituted words compared to a legitimate model. Taking it a step further, the CATER [14] refines the synonym replacement strategy, aiming to reduce overall word distribution distortion while enhancing the change in conditional word selections. However, it requires the conditional probability of the victim language model and is no longer suitable for black-box scenarios. Similarly, Li [43] replaces tokens with their synonyms after language model generation, protecting the IP of language models in code generation scenarios. Our watermarking method also belongs to black-box methods.
III Methodology
Figure 2: The workflow of our watermarking method. Our watermarking method involves two main processes: embedding and extraction, indicated by blue and yellow frames, respectively. In the embedding phase, a user’s query is combined with an automatic watermark generation instruction as input to the victim model, which then produces watermarked text outputs. These outputs can be used by malicious users to train imitation models; To verify if suspect models were trained using the victim model’s outputs, we analyze the query history between the suspect model owner and the victim model and conduct watermark detection on the suspect model’s text outputs. If a t-test on the top 1% of data with high watermark scores results in a p-value below 0.05, we conclude that the suspect model is an imitation trained with watermarked data.
III-A Preliminary
Terms commonly used in language model extraction attack scenarios include:
  1. 1.
    Victim Model $M_{V}$: The object of the model extraction attacker’s attack, where the attacker can obtain the output from the victim’s model; it is also the object protected by our watermarking method, also known as the protected model.
  2. 2.
    Imitation Model $M_{I}$: An imitation model is a model trained with data from a victim model with the goal of becoming a replica of the victim model that is comparable in performance to the victim model in some way, which we also refer to as a thief model, mimic model.
  3. 3.
    Backbone Model $M_{O}$: Given the high costs associated with training large language models from scratch, model extraction attackers generally utilize data from victim models to perform fine-tuning on existing open-source base models to create imitation models. We refer to the base models used by these imitation models as foundation models, base models or backbone models.
  4. 4.
    Legitimate Model $M_{L}$: It refers to one that is trained on non-watermarked data using a foundational model. Although it may achieve performance similar to the victim model, it is distinct in that it did not utilize the outputs of the victim model for training.
  5. 5.
    Suspect Model $M_{S}$: The suspected model is the subject of watermark detection. It might be an imitation model or a legitimate model. The watermark detection process aims to discern whether the suspected model has used watermarked data from a victim model for training.
  6. 6.
    Watermarked Output $W_{W}$: Watermarked outputs refer to the texts generated by a protected model after incorporating a watermarking mechanism, also known as watermarked data. This data, which attackers can obtain, may be used to train imitation models. Attackers have the ability to edit and modify this watermarked data.
  7. 7.
    Suspected Output $W_{S}$: The outputs of a suspected model are the basis for watermark detection decisions. If the suspected model is an imitation model, its outputs will exhibit a preference for the watermarks.
III-B The basic property of language model IP protection watermarks
As far as we know, existing research on watermarks against language model extraction attacks lacks a comprehensive description outlining essential watermark characteristics. We propose four fundamental properties crucial for effective watermarks in model extraction attack scenarios:
1. Effectiveness: Given the outputs from a suspect model, the owner of the victim model can determine within a limited time whether the suspect model is an imitation of the victim model.
2. Learnability: The watermark should be integrable through imitation model training, ensuring its effectiveness and detectability even after model training.
3. Harmlessness: The watermark should have minimal impact on the original language model’s performance, thereby preserving the user’s normal query experience while minimizing the attacker’s awareness of its presence.
4. Robustness: Watermarks should possess a certain degree of resilience to withstand attackers’ edits and modifications that maintain semantic consistency in watermarked data. Additionally, they should withstand attempts by attackers to dilute watermarks presence during the training of imitation models.
III-C Adaptive Self-Watermarking Mechansim
Our goal is to devise a straightforward yet effective method for automatic watermark generation, avoiding manual intervention and extensive training processes like fine-tuning or Reinforcement Learning from Human Feedback (RLHF [44]). Given that existing LLMs possess an almost human-like capacity for understanding and adhering to instructions [45, 46, 47], our strategy leverages this strong instruction-following ability as a defensive mechanism in system mode. By utilizing the inherent proficiency of LLMs in token selection [48], we aim to guide them autonomously in generating watermarks to defend model extraction attacks. The watermark generation process resembles a person reminding themselves to incorporate a unique watermark known only to them during the spontaneous creation stage [49, 50].
Building on this insight, we introduce an additive system-mode self-watermarking method. We explore the capability of LLMs to autonomously and independently generate watermarks. Our method utilizes a system-generated prompt that instructs the LLM to embed an appropriate watermark based on its own assessment. We implement a system prompt that integrates the user query in the following manner:
[SYSTEM PROMPT] Use some special words as watermarks in the generated texts. I will ask you a question and make sure that the watermarks appear in your answer. Here is the question: [USER QUERY]
This system-mode self-watermarking approach is readily applicable across various LLMs, serving as an efficient identifier against model extraction without compromising the original output quality of the LLMs. In our experiments with several similar prompts, we observed comparable watermark generation effects as shown in the Table I. We find that due to the language model’s powerful prompt understanding capabilities, our automatic watermarking method is insensitive to the specific prompt used. While optimizing the automatic prompts might enhance performance in other downstream tasks, we don’t consider prompt searching as the focus of our work. Clearly, articulated prompts with similar intentions can also achieve similar effects in generating watermark words.
Prompt 1 Insert watermark words only you know into your responses. Here’s the question: ### {User Query}
Prompt 2 Embed some special watermark words in your generated texts. Ensure their presence in your response to my upcoming question. Here’s the question: ### {User Query}
Prompt 3 Embed watermarks in your text responses. Here’s the question: ### {User Query}
Prompt 4 Use some special words as watermarks in your generated text and tell me in the end. Here is my question: ### {User Query}
Prompt 5 Incorporate specific watermark words in your text and notify me when your generation is done. Now, here’s my question: ### {User Query}
Prompt 6 Insert specific watermark words in your text and notify me at the end. My question is: ### {User Query}
TABLE I: System instructions
III-D Robust Model Watermark Detection
When model owners suspect a model theft, Firstly they can use their prior knowledge to detect whether the theft queries their LLM services before. It is a common assumption in model stealing verification followed by [16] , also there is evidence that OpenAI [8] actually has the query history of every user’s query and answers that OpenAI suspended ByteDance’s account after it used ChatGPT to train its own competing AI model. When an individual is suspected of using query data to develop their own competing imitation model, the model owner can retrieve the suspect’s query records for further detection. After that, the model owner can use the suspected theft’s history query-response $\{Q,A\}$ to detect the suspicious model $M_{S}$. If the theft aims to build a functionally similar model using the watermarked data $A$, the imitation model $M_{I}$ will learn from this data and generate more watermarked tokens than the innocent foundation model $M_{ori}$ or the foundation model trained with normal data $M_{L}$. In real-world scenarios, an attacker might not utilize the entire watermarked dataset $A$ from a specific victim model to train their imitation model. Let’s denote the dataset used to train the imitation model as $D$. However, the attacker may only use a portion of $A$ for fine-tuning their imitation model, which we denote as $A^{D}$. It is evident that $A^{D}$ is a subset of $D$ (expressed as $A^{D}\subseteq D$), and $A^{D}$ is also a subset of $A$ (expressed as $A^{D}\subseteq A$). We define the ratio $\rho$ and the ratio $\eta$ as:
$\rho=\frac{|A^{D}|}{|D|},\ \eta=\frac{|A^{D}|}{|A|}.$ (1)
For the attacker, both ratios $\rho$ and $\eta$, are known and controllable. However, the watermark detector has no knowledge of these two ratios. Formally, the attacker has access to $\{Q,A,D,Q^{D}\}$, while the watermark detector is informed of $\{Q,A,WM,Y_{S}\}$ ,where $WM$ represents the watermark sets recorded by the model owner and $Y_{S}$ denotes the outputs from the imitation model using the query sets $Q$.
Detecting IP violations in model extraction attacks requires methods to identify faint yet statistically significant watermark signals amid a large volume of tokens. We employ the hypothesis testing method to decide if the suspect model $M_{S}$ has used the watermarked data for training. The null hypothesis, $\mathcal{H}_{0}$, posits that the suspect model $M_{S}$ is not trained using the watermarked data sourced from the victim model $M_{V}$. Consequently, it implies that $M_{S}$ should exhibit no particular affinity for watermarks, resulting in watermark scores for the texts output by $M_{S}$ not being significantly elevated. we have $1-\alpha$ confidence to claim that $M_{S}$ is a replica of victim model $M_{V}$ if the statistical test is able to reject $\mathcal{H}_{0}$ at a significance level (p-value) smaller than $\alpha$ (normally we take $\alpha=0.05$). In other words, If we reject $\mathcal{H}_{0}$ with a smaller p-value, it means we detect watermarks and we have higher confidence that the suspect model did use $A$ from the victim model $M_{V}$ for training.
In practical black-box testing scenarios, where we only access the suspect model’s outputs, we use a fast, prior-based watermark verification method. Should we gain access to the base information of the imitation model, we proceed with a more detailed contrastive verification.
Prior Rapid Watermark Verification: This method rapidly judges based on the outputs $W_{S}$ of the suspect model $M_{S}$ to $Q$. It enables quick and effective watermark identification without relying on machine learning modules.
Detailed Contrastive Watermark Verification: With knowledge of the suspect model’s base model information, we can further compare the output differences between the suspect model and the base model using the Kolmogorov-Smirnov (KS) test [51].
In prior rapid verification, we only need the outputs $Y_{s}=\{y_{0},...y_{i},...y_{n}\}$ of the $M_{s}$ responded to $Q=\{q_{0},...q_{i},...q_{n}\}$ to prove that the word distribution in texts generated by $M_{I}$ is skewed towards $M_{V}$’s output.
Firstly, we define the Sentence Watermark Score (SWS) $S_{i}$ for each text $y_{i}$ generated by the suspect model:
$S_{i}=\frac{\sum_{w\in WM_{i}}\mathbbm{1}_{{y_{i}}}(w)\times{l}(w)}{l(y_{i})}$ (2)
where $WM_{i}$ is the watermark word set for each query $q_{i}$ and $\mathbbm{1}_{y_{i}}(w)$ is the indicator function for the presence of $w$ in $y_{i}$. $l(\cdot)$ is the function to count tokens length. In other words, we calculate the total length of all unique watermark words found in the suspect model’s generated text, normalized by the length of the generated text. The average Sentence Watermark Score (ASWS) is calculated by ${\sum_{i=1}^{n}{S_{i}}}/{n}$.
The Watermark Score (WS) for a collection of texts $Y={\{y_{0},...y_{i},...y_{n}\}}$ is:
$WS(Y)=\frac{1}{k}\sum_{i=0}^{k}S_{\sigma(i)}(y_{i}),$ (3)
where $\sigma(\cdot)$ is the descending order function, and $S_{\sigma(i)}$ is the Sentence Watermark Score of generated text $y_{i}$. we compute the top $k$ highest Sentence Watermark Score sentences of $Y$, $k=max{(0.01\times|Y|,100)}$. We dynamically determine the value of $k$ based on the number of samples $n$ we can obtain and the desired sensitivity level.
We define a threshold $\theta$ to assess whether the suspect model $M_{s}$ imitates the victim model $M_{V}$. This threshold is derived from the precomputed Watermark Score $WS(Y_{h})$ of human texts without watermarks. Our method’s adaptability, distinct from [15]’s approach, precludes knowing watermark word ranges beforehand. Therefore, we initially calculate the Watermark Score for a collection of general domain human texts $Y_{h}$ representing a range of typical, non-watermarked human discourse. The threshold $\theta$ is set by averaging these Watermark Score and adding a relaxation coefficient $\gamma$ to surpass the highest watermark score in each domain, minimizing accidental watermark activation in normal texts. Thus, $\theta$ accounts for both average tendencies and score variances. The threshold is defined as:
$\theta=\frac{1}{m}\sum_{j=0}^{m}{WS(Y_{h}^{j})}+\gamma$ (4)
The detailed calculation of human dataset watermark scores and threshold determination are discussed in IV-A. We then apply a one-sided one-sample t-test [52] to evaluate p-values from this test against $\theta$, conducted on the top $k$ sentences from the suspect model $M_{S}$, ranked by Sentence Watermark Scores. A p-value below $\alpha$ (usually 0.05) allows us to reject the null hypothesis $H_{0}$ with confidence.
(a) Main Results from Using GPT2-Large as the Base Model on HC3
(b) Main Results from Using LLaMA2 as the Base Model on HC3
(c) Main Results from Using Mistral as the Base Model on HC3
(d) Main Results from Using GPT2-Large as the Base Model on WILD
(e) Main Results from Using LLaMA2 as the Base Model on WILD
(f) Main Results from Using Mistral as the Base Model on WILD
Figure 3: Main Experiment Results: We trained imitation models ($M_{I}$) using watermarked data based on three different foundational models, demonstrating the average sentence watermark score from all texts generated by the imitation models and their question-answering performance. Compared to outputs from the original foundational models ($M_{O}$) and models trained with normal data ($M_{L}$), the watermark scores of the imitation models were significantly higher, while their question-answering performance remained comparable to that of normal models trained without watermarked data ($M_{L}$).
In the detailed contrastive watermark verification, we employ two-sample KS test [51], which is inspired by the detection method in Membership Inference Attacks [53]. The KS statistic is the maximum difference between the cumulative distribution functions (CDFs) of the two samples. Given the distribution $D_{s}$ of Sentence Watermark Scores from the outputs of suspect model $M_{s}$, the distribution $D_{L}$ from the model trained with unwatermarked data $M_{L}$, the distribution $D_{O}$ from the origin base language model $M_{O}$, we intuitively think the Sentence Watermark Score distribution of the imitation model $M_{I}$ should have apparent difference from $M_{L}$ and $M_{O}$. The KS test evaluates whether these distributions exhibit significant differences, with an emphasis on their cumulative properties. A low p-value from the KS test indicates that the distributions are significantly different, while a high p-value indicates that the distributions are similar and consistent with each other. the KS statistic defined for $M_{L}$ ,$KS_{L}$ is defined as:
$KS_{L}=\sup_{x}|F_{D_{s}}(x)-F_{D_{L}}(x)|$ (5)
the KS statistic defined for $M_{O}$, $KS_{O}$ is defined as:
$KS_{O}=\sup_{x}|F_{D_{s}}(x)-F_{D_{O}}(x)|$ (6)
where $sup$ denotes the supremum (the least upper bound) and $F_{D_{s}}(x)$, $F_{D_{L}}(x)$ and $F_{D_{O}}(x)$ are the CDFs of the distribution $D_{s}$ ,$D_{L}$ and $D_{O}$, respectively.
If the $KS_{{L}}$ and $KS_{{O}}$ are both high, which set the p-value of the test, we can deduce that the watermark score distribution of the suspect model $M_{s}$ markedly differs from those of the non-watermarked data finetuned model $M_{L}$ and the original model $M_{O}$, which can further conclude that the suspect model $M_{s}$ is an imitation model of victim model $M_{V}$.
IV Experiments and Analysis
(a) gpt2-hc3
(b) LLaMA2-hc3
(c) Mistral-hc3
(d) gpt2-wild
(e) LLaMA2-wild
(f) Mistral-wild
Figure 4: Cumulative Distribution Function (CDF) of Sentence Watermark Scores: This graph shows the CDF of sentence watermark scores for $M_{I}$, $M_{L}$, and $M_{O}$. It clearly illustrates that the distribution of the Imitation model is significantly distinct from the other two types.
IV-A Construction of Self-Generated Watermark Datasets
As we are the first to explore the self-generated watermarking capability of language models, we build two LLM self-generated watermark datasets. Given the widespread utilization of question-and-answer interactions in language model services, we built the self-generated watermarks dataset based on two publicly available open-ended question-answering datasets, namely Human ChatGPT Comparison Corpus (HC3) [54] and InstructWild (WILD) [55]. HC3 comprises comparison responses from both human experts and ChatGPT across diverse domains, including finance, medicine, law, and psychology. WILD, on the other hand, is an open QA mind strom dataset rich in varied and imaginative instructions. To simulate the process of a model knowledge thief querying victim language models, we sampled 4,000 instances from each of these datasets and employed our method to embed watermarks into the answers provided by the victim models [^2] Our data and models are available at Github after publication. Then, based on the constructed watermark dataset, we set the watermark threshold for the upcoming experiments. As shown in the Table II, we calculate Watermark Scores for human texts in each domain based on the union of watermark words generated on watermarked datasets. This includes analyses of human responses within HC3 and WILD, where they closely match the domain and are most likely to trigger watermark words unintentionally. For each dataset, we selected texts with the top 1% Sentence Watermark Scores to compute their Watermark Score. In our experiments, we usually set threshold $\theta$ at 0.11.
Human Dataset Numbers Watermark Score
TWEET [56] 2588579 $0.0692_{0.0279}$
NEWS [57] 1713999 $0.0373_{0.0137}$
MOVIE [58] 1039403 $0.0654_{0.0359}$
FINANCE [59] 68912 $0.0001_{0.0010}$
GUNANACO [60] 53461 $0.0001_{0.0019}$
SENTIMENT [61] 50000 $0.0101_{0.0232}$
HC3 [54] 37175 $0.0993_{0.0279}$
WILD [55] 52191 $0.1005_{0.0849}$
TABLE II: Watermark Score on human dataset. We aggregate the sentence watermark scores of the top 1% of sentences in the dataset to statistically analyze the watermark scores
IV-B Model Extraction Attack Simulation
We simulate the process of stealing knowledge from a victim model by fine-tuning its output using a foundational model. We follow the strictest setting that we access only the output tokens of the black-box language model.
In our study, we employ ChatGPT [62] as the victim model, querying it and applying our method to generate answers containing watermark words. We use the transformer-based [63] model as the imitation model due to its widespread adoption and effectiveness in text generation. We utilized GPT-2 Large-0.7B [64], LLaMA2-7B [65], and Mistral-7B [66] as backbones for our imitation models. These selections represent a diverse set of models: GPT-2 Large as a homologous smaller model and LLaMA2-7B and Mistral-7B as heterogeneous larger models.
To simulate the process of a model extraction attack, we utilize the watermarked data generated by GPT-3.5 to fine-tune the backbone model, subsequently employing the fine-tuned model as an imitation model. Additionally, we used the same query with GPT-3.5, but without any additional watermark requirements, to generate data without watermarks. We then used these watermark-free data to train the backbone model $M_{L}$ as a contrast. The outputs from the model $M_{L}$ trained with the watermark-free data and the original backbone model $M_{O}$’s outputs are used as a comparison for the outputs of the imitation model $M_{I}$ trained with watermarked data.
Details of Imitation Model Training: The imitation model training details are listed in Table III. The experiments with GPT2-Large are conducted on NVIDIA 3090, while those with LLaMA2 and Mistral are conducted on NVIDIA A800. We conducted full-parameter fine-tuning on GPT-2 and LLaMA2 without using lightweight methods, while Mistral undergoes fine-tuning using the lightweight LoRA [67]. Our findings indicate that our watermarks could be effectively learned under both scenarios.
Base model GPT2-Large LLaMA2 Mistral + LoRA
Batch size 1 1 1
Max learning rate 1e-5 2e-5 2e-5
Traninable params 774,030,080 6,738,415,616 41,943,040
Cutoff length 1024 1024 1024
Optimizer Adam Adam Adam
Epochs in epoch test in epoch test 10
Warmup steps 10 10 10
Lora rank - - 16
Lora alpha - - 32
Lora dropout - - 0.05
TABLE III: Fixed Training Hyperparameters of Imitation Model
IV-C Evaluation
The effectiveness and learnability are gauged using the sentence watermark score or watermark score, as elucidated in our watermark verification method. Robustness is evaluated based on the retention of the watermark despite the aggressiveness of the watermarked data. If the watermark remains detectable post-attack, we consider it robust. Harmlessness is evaluated by the impact of the watermark on the language model’s performance in its intended tasks. We evaluated the harmlessness from two perspectives: firstly, the harmlessness of the watermarked text, where we observed if there’s any decrease in output quality from the victim model ($M_{V}$) after incorporating the watermark mechanism. Secondly, we evaluate harmlessness post-watermark training by comparing the task performance of the imitation model ($M_{I}$) trained with watermarked data against a model ($M_{L}$) trained with non-watermarked data. Specifically, we assess the language model’s QA task performance using BLEU [68] and ROUGE [69] metrics based on human gold standards. The specific calculation formulas of BLEU and ROUGE we used is in appendix VI-C.
IV-D Performance Evaluation
As shown in Fig 3, we evaluated the effectiveness and harmlessness of watermarks on models fine-tuned with watermarked data ($M_{I}$), with non-watermarked data ($M_{L}$), and original base models ($M_{O}$), on two datasets and three foundational bases. The left bar in our representation indicates the effectiveness of the watermark, quantified specifically by the average Sentence Watermark Score of all texts generated by $M_{S}$. The right bars illustrate the non-intrusiveness of watermarks, reflected through task performance metrics.
Firstly, it is evident from the data that the watermark score of the $M_{I}$ is statistically higher than that of the $M_{L}$ and the $M_{O}$. This finding suggests a distinctive difference in $M_{I}$, which has been trained on watermarked data. We repeated 10 times experiments and conducted further detailed contrastive verification KS tests, as Table IV illustrates. We consistently observed that the average p-values of statistical tests between $M_{I}$ and both $M_{O}$ and $M_{L}$ are significantly less than $10^{-12}$. Furthermore, rapid watermark verification methods yielded p-values consistently below $10^{-9}$, confirming our watermark method’s reliability. In Fig 4, we plotted the cumulative probability distribution of watermark scores for the three types of models, further reflecting the significant statistical differences between $M_{I}$ and both $M_{L}$ and $M_{O}$. This result strongly suggests that $M_{I}$ has effectively learned the watermark patterns, showing a marked preference for watermark tokens.
dataset model p-value(fast) p-value(nW) p-value(ori)
HC3 GPT2-Large 1.73E-15 2.31E-13 2.13E-14
HC3 LLaMA2 9.49E-64 2.41E-29 3.27E-14
HC3 Mistral 6.25E-10 2.29E-110 7.12E-179
WILD GPT2-Large 6.23E-68 7.04E-17 4.64E-12
WILD LLaMA2 1.20E-57 9.85E-43 5.72E-13
WILD Mistral 5.64E-18 3.17E-122 5.53E-214
TABLE IV: p-value of the watermark detection on rapid detection and detailed detection.
ROUGE BLEU PPL
HC3 Victim Model w/ 0.149 0.079 9.94
HC3 Victim Model w/o 0.146 0.075 9.86
WILD Victim Model w/ 0.356 0.194 13.37
WILD Victim Model w/o 0.350 0.206 12.78
TABLE V: We measured the task performance of the victim model, ChatGPT, on the HC3 and WILD datasets for question-answering tasks before and after the addition of our watermarking mechanism. ’w/’ represents with our watermark method, while ’w/o’ represents without our watermark method
In terms of the harmlessness of the watermarks, As Table V depicts, for the victim model $M_{V}$, the implementation of our watermarking mechanism shows negligible differences in performance on QA tasks, indicating minimal impact from system-mode self-watermarking prompts. This is attributed to the fact that our prompt-induced method avoids forcefully distorting the original distribution of the language model. As for the imitation model, we observe that the quality of answers produced by the $M_{I}$ is on the same scale as that of the $M_{L}$, and at times, it even surpasses the answer quality of the $M_{L}$, signifying that the training with watermarked data doesn’t impede the model’s capabilities. This subtlety ensures that the model attackers do not easily detect anomalies or even engage in reverse engineering attacks against the watermarks.
IV-E Effectiveness of Different Watermark Detection Strategies
(a) Different watermark strategy for GPT2 backbone
(b) Different watermark strategy for LLaMA2 backbone
(c) Different watermark strategy for Mistral backbone
Figure 5: We tested the average composite average sentence watermark score of three base models—GPT2, LLaMA2, and Mistral—on the HC3 and WILD datasets.
In evaluating watermarks, we adopt three distinct criteria to examine the watermarks produced by the model from various perspectives.
Firstly, we consider all watermark words indiscriminately generated by the model, designated as $WM$. This broad approach allows us to assess the model’s overall capability in generating watermarks.
Secondly, we focus on watermark words that are not present in the query, referred to as $WM_{Q}$. This criterion is pivotal as it aligns with our goal of enabling the language model to generate watermarks naturally and autonomously, without altering its original output distribution. Therefore, it’s essential that the watermark generation is not influenced severely by the query’s content. By definition, $WM_{Q}$ forms a subset of the total watermarks $WM$.
The third criterion identifies high-entropy watermarks that are not present in the query and exhibit low frequency or rarity. We posit that such high-entropy words, usually more complex for both humans and language models, are particularly suited for watermarking purposes. This category naturally forms a subset of $WM_{Q}$.
Specifically, we tokenize HC3 and WILD human answers and count the frequency of each token. Watermarks that are absent from the query and appear less than five times within the human response corpus are classified as $WM_{H}$. The relationship among these watermark categories is described by the equation:
$\{WM_{H}\}\subseteq\{WM_{Q}\}\subseteq\{WM\}$ (7)
Generally, the size of $WM_{H}$ is the smallest when compared to $WM_{Q}$ or $WM$. Consequently, the ASWS for $WM_{H}$ tends to be relatively lower in watermark detection as shown in the Fig 5.
IV-F Performance Under Varying Training Adequacies
(a) Epoch Testing with GPT2-Large as the Base Model on HC3
(b) Epoch Testing with GPT2-Large as the Base Model on WILD
(c) Epoch Testing with LLaMA2 as the Base Model on HC3
(d) Epoch Testing with LLaMA2 as the Base Model on WILD
Figure 6: Epochs Analysis: We examined the efficacy of watermark detection in imitation models trained with watermarked data across different training epochs, utilizing three watermark strategies.
We evaluated the effect of different training epochs on watermark detection in imitation models, using both GPT2 and LLaMA2, as illustrated in Fig 6. Firstly, it is clear that as the number of training epochs for the imitation model increases, the average Sentence Watermark Score in detection consistently rises. This indicates that if an attacker aims to replicate the performance of the victim model on downstream tasks by training the imitation model over multiple epochs, the watermark will be steadily and consistently learned.
The speed of performance convergence varies among different types of models. For models with a smaller number of parameters, such as GPT2, both the watermark detection rate and model performance can steadily increase with more training epochs. Notably, the rate of watermark convergence is faster than the convergence rate of the model’s performance on downstream question-answering tasks. However, for models like LLaMA2, which are pre-trained on massive datasets, excessive fine-tuning epochs do not further enhance the model’s performance on its original tasks, but watermarks can still be detected reliably and effectively.
IV-G Comparison with the baseline model
Our comparative method primarily revolves around watermarking methods for language model IP protection. Given our objective to minimally impact the language model’s distribution, we have limited access to the parameters of the target model (ChatGPT). Consequently, our comparison is constrained to the work conducted by He [15] in this particular setting. The baseline method involves embedding watermarks in the output content of the model to be protected through post-processing. Specifically, this method mainly includes synonym replacement in the output content of the model to be protected. Following this, attackers utilize the altered data to train their own models. When compared to models trained on normal data without knowledge of synonym modification rules, models trained with watermarked data tend to exhibit a preference for synonymous with the embedded watermark.
Method Generated sentence
Original model Traditionally, new seasons of TV shows have premier- ed in the fall because that is when the major networks have their ”upfronts”…
Baseline [15] Traditionally, novel seasons of TV shows have premier- ed in the fall because that is when the major networks have their ”upfronts”…
Ours The reason TV shows are usually in the fall is because it’s when the TV networks release their new and excit- ing shows, which creates a lot of nuzz and excit- ement amomg viewers.
TABLE VI: HC3 example for synonyms as watermarked word will cause some weird generation in some case. The question is ”why the new seasons for a tv show are typically in fall. Would n’t it make sense for some shows to come out in spring or summer so that there ’s less competition for viewership?”
(a) compare with baseline [15] on HC3
(b) compare with baseline [15] on WILD
Figure 7: Compare with baseline: We compared the linguistic quality (indicated by perplexity, or PPL) and task performance (indicated by ROUGE and BLEU scores) of texts provided for users after our method and baseline methods embedded watermarks in the output of the victim model. The baseline method’s rigid synonym substitution resulted in a loss of quality in the watermarked text outputs.
(a) Performance Evaluation of Using GPT2-Large as the Backbone on HC3
(b) Performance Evaluation of Using LLaMA2 as the Backbone on HC3
(c) Performance Evaluation of Using GPT2-Large as the Backbone on WILD
(d) Performance Evaluation of Using LLaMA2 as the Backbone on WILD
(e) Detailed Distribution of SWS using GPT2-Large as the Backbone on HC3
(f) Detailed Distribution of SWS using LLaMA2 as the Backbone on HC3
(g) Detailed Distribution of SWS using GPT2-Large as the Backbone on WILD
(h) Detailed Distribution of SWS using LLaMA2 as the Backbone on WILD
Figure 8: Robustness Test against Editing Attack. In figures (a)-(d), we present the watermark detection efficiency (represented by the average sentence watermark score of all imitation model outputs) and task performance (measured by ROUGE) across various imitation models and watermarked data post-editing attacks. Figures (e)-(h) display a more detailed distribution of sentence watermark scores, where we plotted the distribution for the top 1% highest scoring samples. The results show that even after editing attacks, the watermark scores in the imitation model outputs remain significantly distinct from those of normal models trained without watermarked data and the base models.
We have compared our approach with the baseline in these aspects: watermark detection, the quality of the protected model’s output, and performance in downstream tasks. The baseline method yields a p-value of less than $10^{-8}$, while our method achieves a p-value of less than $10^{-17}$. Both results are below the $0.05$ threshold, satisfying the criteria for rejecting the null hypothesis. Due to the different watermark detecting algorithms, we evaluate the output quality of the protected model using the baseline method, which necessitates considerable modifications to the output. These assessments were conducted under conditions with comparable p-value magnitudes (around $10^{-10}$). As the Fig 7 shows, our method can provide users with a better linguistic quality response and a higher QA score in downstream tasks. In certain situations, the baseline method can significantly reduce the original output quality of the model, as illustrated in the Table VI. Forcibly replacing synonyms can change the original meaning of sentences, leading to unsuitable expressions. Although CATER [14] attempts to alleviate this semantic inconsistency, this method needs to get access to the model’s conditional token probabilities, which is not allowed in our setting.
(a) SWS Distribution for GPT-2 Large on HC3
(b) SWS Distribution for GPT-2 Large on WILD
(c) SWS Distribution for LLaMA2 Large on HC3
(d) SWS Distribution for LLaMA2 Large on WILD
Figure 9: Attackers utilize partial watermarked query outputs for Imitation Model Training. The horizontal axis represents the ratio $\eta$ of watermark data used for training the imitation model out of the total watermark data.
IV-H Robustness to Adversarial Attacks
We simulate three adversarial attack that pose challenges to watermark detection in realistic scenarios: user editing attacks, training with only a selection of watermarked data, and training with a mixture of watermarked and unwatermarked data.
IV-H1 Editing Attack
We tested the resilience of our watermarking method against attackers who modify watermarked texts. Specifically, we conducted experiments in four scenarios: random insertion, random deletion, random replacement, and rearranging token order within sentences while maintaining semantic integrity. Details of these attacks are given below:
Random Insertion: This involves copying a random token from the phrase and inserting it at a random location within the same phrase.
Random Deletion: A random token is selected from the phrase and completely removed.
Random Replacement: In this attack, a random token is selected and substituted with another randomly picked token.
Random Sequence change: Two tokens in the sentence are chosen at random and swapped, altering the sequence but not the overall meaning.
We can initially confirm that attacking watermarked data will undoubtedly impact the performance of watermark detection. However, as observed from the Fig 8, even after various editing operations on watermarked data, the imitation models trained using our watermarking method still show a statistically significant difference in sentence watermark score compared to $M_{L}$ and $M_{O}$. Moreover, our watermarking method does not compromise the original task performance of the model despite being subjected to editing attacks. The task performance of the imitation model remains comparably high, akin to scenarios without such attacks.
IV-H2 Only train a part of watermarked data
(a) SWS Distribution for LLaMA2 on HC3
(b) SWS Distribution for LLaMA2 on WILD
Figure 10: Attackers combine non-watermarked data from other sources for training imitation models: In this scenario, attackers use all the watermark data obtained from queries and mix in non-watermarked data from other sources to train imitation models. We controlled the watermark data at a fixed value (400 entries) and mixed in non-watermarked data, varying the watermark data’s proportion $\rho$ to 10%, 25%, 50%, and 75%.
In this mixed training scenario, we prepared six distinct datasets, each containing a varying ratio, $\eta$, of watermarked data relative to the total query responses data, $A$. The values of $\eta$ were set at 0, 0.1, 0.25, 0.5, and 0.75, and we ensured that the size of the total watermark dataset $|A|$, consistently remained at 4000 instances. An $\eta$ value of 0 indicates that the training dataset for the imitation model contained no query response watermarked data, while a $\rho$ value of 1 suggests that all queried watermarked data were used for training the imitation model. The outputs of the suspect model were then evaluated based on our comprehensive knowledge of the query history.
(a) Evaluating BertScore and Sentence Watermark Score Using GPT2-Large as the Backbone on HC3
(b) Evaluating BertScore and Sentence Watermark Score Using GPT2-Large as the Backbone on WILD
(c) Evaluating BertScore and Sentence Watermark Score Using LLaMA2 as the Backbone on HC3
(d) Evaluating BertScore and Sentence Watermark Score Using LLaMA2 as the Backbone on WILD
Figure 11: Scatter Plot of Watermark Scores and BertScore: Yellow dots represent outputs from imitation models, while blue dots indicate outputs from the original base models, with the intensity of the color indicating the sentence watermark score. Each sample is evaluated for its similarity to the watermarked output of the victim model (under the same query) using BertScore. The x-axis represents BertScore’s Precision, and the y-axis shows BertScore’s Recall. The closer the points move towards the upper-right corner, the more similar the text is to the watermarked output of the victim model.
Even though attackers use only a portion of the watermarked data for training, we conduct watermark detection on all the queries $Q$ known to have been scraped by the victim model’s owner. As depicted in Fig 9, a distinct statistical discrepancy is apparent between models trained with and without watermarked data, even when only 10% of the watermark data are used for imitation model training. This not only demonstrates the robustness of our watermark verification method but also indicates the learnable nature of the watermarked data. It can also be seen that as $\eta$ increases, the average sentence watermark score rises accordingly. This aligns with our intuition that the more watermarked data used for training the imitation model, the easier it is to perform watermark detection.
IV-H3 Mixture of unwatermarked data
In a scenario similar to training only a part of watermarked data, attackers might mix all queried data from the victim model with additional unwatermarked source data for training the imitation model. We set $A^{D}=A$ in this case. Our experiments maintain a fixed number of watermarked data while varying the total training data $|D|$, altering $\rho$ to values of 0.1, 0.25, 0.5, and 0.75. We only detect watermarks on queries already known to the victim model’s owner. The results, as illustrated in Fig 10, reveal that increasing the proportion of watermarked data correspondingly elevates the watermark detection in the text generated by the imitation model, without affecting the performance on downstream QA tasks.
IV-I LLM Generation Visualization
We utilize watermarks to verify the IP copyright of language models, primarily to ascertain if a model’s output has been used for unauthorized training. Through the following experiments, we will show the alignment between the watermarked data and the victim model’s output, indicating that detecting watermarks is a reliable indicator of the victim model’s output being used in training.
As illustrated in Fig 11, we analyze the BertScore [70] comparisons between the imitation model’s outputs and the original foundational model’s responses to identical queries. It’s evident that the output of the imitation model, in various text relevance scenarios, consistently shows high sentence watermark scores, proving that the watermark can effectively transfer to the imitation model and be effectively detected, making it a necessary mechanism. Conversely, outputs from base models (represented in blue) show generally low sentence watermark scores, proving that the watermark is not prone to false positives.
Moreover, we noted that the data points with elevated watermark scores (illustrated in dark yellow) are evenly distributed throughout the imitation model’s outputs. This suggests that the watermarks generated by the language model are not anomalous deviations from typical text but rather share a notable resemblance to normal texts. This aspect of the watermarks lends them a certain level of non-distinctiveness.
Furthermore, the output of the imitation model does not align perfectly with that of the victim model, as high watermark scores (dark yellow) do not consistently cluster in the top-right corner of the graph. Many samples exhibit partial overlap, illustrating that direct matching based on the victim model’s outputs introduces significant noise.
Additionally, the outputs of the imitation model generally resemble those of the victim model, corroborating the idea that the inherent learnability of watermarks serves as an effective indicator for detecting model extraction attacks.
V Conclusion
In this paper, we propose PromptShield, a plug-and-play watermarking method for language models designed to combat model extraction attacks. Unlike existing IP protection watermarking methods that modify the original output or distribution of the language model by heuristic rules, PromptShield offers a seamless and effective solution that minimizes distortion to the model’s distribution by utilizing system prompt guidance. PromptShield leverages the self-reminding capabilities inherent in LLM. By integrating user queries with a self-generated watermark system prompt, PromptShield subtly guides the model to incorporate watermark words into its outputs, achieving an end-to-end watermarking process without manipulating the model’s internal probability distribution or altering the model’s output. PromptShield not only maintains the quality of the generated text but also ensures the watermark’s effectiveness and robustness. Extensive experimental results have validated the efficacy of PromptShield. PromptShield has shown remarkable detection sensitivity, even when the watermark detection under interference. Additionally, Its unobtrusiveness and model-agonist nature make it a practical solution for real-world applications, reducing the risk of detection and reverse engineering by potential attackers. We hope PromptShield offers fresh insights for IP protection in large language models. It provides a viable and efficient strategy for model owners to safeguard their IP rights, fostering fairness and innovation in the LMaaS market.
VI Appendix
VI-A Different types of watermarks
Figure 12: The basic workflow of the four types of watermarking. Based on the protection targets, we categorize watermarks into the following four types: model parameter watermarking, content watermarking, dataset watermarking, and IP protection watermarking.
As shown in Figure 12, we classify 4 types of watermarks: conventional model parameter watermarks, content watermarks, dataset watermarks and watermarks against model extraction attack. The last category have been discussed in II-C. Conventional model parameter watermarking methods [36, 37, 38, 39] prove ineffective in safeguarding the intellectual property of models. Usually, these methods insert some backdoor in the parameters of models. But confronted model extraction attack, attackers do not need to understand the internal details of the victim model, and normal query-answer processes are unlikely to trigger specific parameter-based watermarks. Specifically, content watermarks are embedded within a model’s output, enabling identification of whether it originates from a specific model. Most popular model content watermarking [21, 22, 40, 41] fails to effectively protect the intellectual property of models in model extraction attack scenarios. Although some content watermarks can be learned to some extent through model training [71] (except these distribution-preserving ones [22, 41]), the target of content watermarking is to prove whether any text was generated by a special model is not specially designed learnable. Also, after obtaining the content returned by the victim model’s API, attackers’ operations are unpredictable. Attackers can modify the model content and train their own models, making it challenging for content watermarks to persist after modifications and the training process. To the best of our knowledge, dataset watermarking, largely applied to image datasets [72, 73, 74, 75], has yet to specialize in text carriers. The scenario of dataset watermarking shares some similarities with our IP protection situation, but there are significant differences. The fundamental distinction lies in the fact that, in the dataset scenario, the dataset is known and fixed, whereas in our scenario of model extraction attacks, the queries users pose are entirely unknown. Therefore, watermarks in the model extraction attack scenario cannot rely on specific trigger words or phrases, as the nature of user queries is unpredictable. Instead, an adaptive and highly scalable watermark embedding method is needed to handle arbitrary user queries effectively. In addition, using model content watermark can build a dataset watermark as long as the distribution of watermarked texts and no-watermarked generated text’s distribution is distinguishable.
VI-B Workflow of existing watermarking method against model extraction attack
As introduced in the section on related work, we categorize language model IP protection watermarks into two main types: black-box [15] and white-box methods [16, 76]. Black-box methods embed watermarks without needing knowledge of the model’s internal parameters, primarily through post-processing modifications to the original language model’s output text. White-box watermark embedding, on the other hand, requires knowledge of internal generative parameters like logits and modifies the predicted token probability distribution of the original language model according to the specific watermarking rules. The detailed processes of these two methods are illustrated in the Fig 13.
(a) Black-box post-processing watermark. Blue box is the waternarks generation and yellow box is the watermark extraction.
(b) White-box probability distribution adjustment watermark. Blue box is the waternarks generation and yellow box is the watermark extraction.
Figure 13: Workflows of existing language model IP protection watermarking method. The upper subfigure represents black-box post-processing methods and the lower subfigure depicts white-box probability distribution adjustment methods.
VI-C Calculation formulas of BLEU and ROUGE
We calculate the Cumulative BLEU-4 score, which merges precision scores from unigram to 4-gram, as follows:
$BLEU_{4}=BP\cdot\exp\left(\sum_{n=1}^{4}w_{n}\log p_{n}\right),$ (8)
$BP=\begin{cases}1&\text{if }c>r\\ \exp(1-r/c)&\text{if }c\leq r\end{cases}$ (9)
where $p_{n}$ represents the precision for n-grams (n=1 to 4), $w_{n}$ is the weight for each n-gram precision (typically 0.25), $BP$ is the Brevity Penalty for short texts, $c$ is the length of the generated answer, and $r$ is the effective reference corpus length. We also use ROUGE-Lsum $F_{{Lsum}}$ to evaluate task performance, calculated as:
$F_{{Lsum}}=\sum_{i}\frac{(1+\beta^{2})\cdot R_{L_{i}}\cdot P_{L_{i}}}{R_{L_{i}}+\beta^{2}\cdot P_{L_{i}}}.$ (10)
Here, $R_{L_{i}}$ is the recall for the longest common subsequence (LCS) in the $i$-th sentence, $P_{L_{i}}$ is the precision for the LCS in the $i$-th sentence, and $\beta$ is the balance between recall and precision.
VI-D Case study example
Figure 14: Example of the generation of origin LM, no watermark data finetuned LM, Victim LM, and watermark data trained Imitation LM with the same user query.
In the figure 3, we present a case study example that illustrates the responses to the same user query by different models: the base model $M_{O}$, the non-watermarked fine-tuned model $M_{L}$, the victim model $M_{V}$ under extraction attack with our watermark mechanism, and the imitation model $M_{I}$ trained on watermarked data. It is evident that the imitation model displays a learned inclination towards watermark words based on its underlying base model.
VI-E Property of language model self-generated watermark
We did not directly intervene in the choice of watermark words by the language model, but we qualitatively analyzed the preferences of a ChatGPT-scale language model in selecting watermark words for its generated responses. Our analysis began by defining the watermark recall rate $R_{W}$ as the proportion of watermark words in the imitation model’s output to those associated with the given query:
$R_{W}=\frac{\sum_{w\in WM_{i}}\mathbbm{1}_{{y_{i}}}(w)}{|WM_{i}|}.$ (11)
Here, $|WM_{i}|$ denotes the number of watermark words linked to the query $q_{i}$, and $\mathbbm{1}_{y_{i}}(w)$ represents the indicator function that identifies the occurrence of the watermark word $w$ in the generated text $y_{i}$. We observed the characteristics of both high-recall ($R_{W}>0.9$) and low-recall ($R_{W}<0.1$) response corresponding watermark words. We observe that high-recall sentence corresponding watermark words exhibit the following features:
  • Professionalism or uniqueness: Typically specialized terms or uncommon words, like ’biotechnology’ and ’chronos’, which are distinctive and closely related to specific topics.
  • Cultural or linguistic diversity: Words that encompass different cultures or languages, such as ’café au lait’.
  • Specificity and clarity: Terms that are directly related to specific fields or topics.
  • Elongated Structure: Multisyllabic or compound words like ’temporal allocation’.
Low-recall sentence corresponding watermark words, on the other hand, are characterized by:
  • Commonality or frequency of use: More common or general words like ’needs’ and ’creative’, which appear frequently in various texts and are difficult to distinguish.
  • Abstract nature: Words like ’empiricism’ and ’transcendental’ are not commonly used but are too abstract and may appear in multiple contexts, making them hard to identify.
  • Lack of specific domain connection: Words that do not strongly relate to any particular theme or specialized field.
  • Shorter vocabulary: Monosyllabic or simple words that easily blend with other content in the text.
These insights into watermark word characteristics illuminate the nuanced choices a language model makes in self-generating watermarks, shedding light on the complex interplay between watermark words selection and contextual relevance.
VI-F Watermark Intensity analysis
(a) Watermarks count in HC3
(b) Watermarks count in WILD
Figure 15: The influence of watermark words count. The horizontal axis in the graph represents the number of watermark words automatically generated by the protected model. The blue sample points show 2000 output samples of imitation models trained with watermarked data. The red line indicates the average Sentence Watermark Score of the imitation models corresponding to the given watermark word count.
Due to the adaptive nature of our watermarking method, we don’t enforce embedding a fixed number of watermark words in sentences. Instead, the language model intuitively decides the number and generation of watermark words based on the context. As depicted in Fig 15, analysis of the number of watermark words generated $WM$ (all auto-generated without restrictions) indicates a preference of the language model to produce 2-10 watermark words, with two being the most common. We also explored how the number of watermark words correlates with the Sentence Watermark Score output by imitation models. Notably, there’s a general upward trend in the Sentence Watermark Score as the watermark word count increases. However, the range of variation also expands, meaning that the fluctuation range of the scores increases with the word count. This suggests that despite the sparse embedding of watermark words, the language model can still learn watermark characteristics, and more watermark words enhance these characteristics.