### v2

# ModelShield: Adaptive and Robust Watermark against Model Extraction Attack

Abstract
Large language models (LLMs) demonstrate general intelligence across a variety of machine learning tasks, thereby enhancing the commercial value of their intellectual property (IP). To protect this IP, model owners typically allow user access only in a black-box manner, however, adversaries can still utilize model extraction attacks to steal the model intelligence encoded in model generation. Watermarking technology offers a promising solution for defending against such attacks by embedding unique identifiers into the model-generated content. However, existing watermarking methods often compromise the quality of generated content due to heuristic alterations and lack robust mechanisms to counteract adversarial strategies, thus limiting their practicality in real-world scenarios. In this paper, we introduce an adaptive and robust watermarking method (named ModelShield) to protect the IP of LLMs. Our method incorporates a self-watermarking mechanism that allows LLMs to autonomously insert watermarks into their generated content to avoid the degradation of model content. We also propose a robust watermark detection mechanism capable of effectively identifying watermark signals under the interference of varying adversarial strategies. Besides, ModelShield is a plug-and-play method that does not require additional model training, enhancing its applicability in LLM deployments. Extensive evaluations on two real-world datasets and three LLMs demonstrate that our method surpasses existing methods in terms of defense effectiveness and robustness while significantly reducing the degradation of watermarking on the model-generated content.
Index Terms:
Large Language Models, Model Extraction Attack, Text Watermarking, Model IP Protection
I Introduction
Figure 1: Language models offering services are at risk of model extraction attacks, which unfold in three steps, depicted in the diagram with blue boxes. Attackers first collect data output by the victim model through queries, then use these data to train their own imitation models. Eventually, the imitation model achieves performance comparable to the victim model, significantly endangering the IP of the victim model. IP protection watermarking involves embedding a special watermark signal in the output of the victim model. When watermarked data are used to train an imitation model, the watermark signal can still be detected in the model’s output. The stages of watermark generation and extraction are highlighted in yellow boxes in the diagram.
In the current era of Language Model as a Service (LMaaS), safeguarding the intellectual property (IP) of large language models (LLMs) becomes a critical issue. The training of LLMs demands significant intellectual efforts, involving massive amounts of high-quality data, extensive computational resources, and human-elaborated training design. These investments imbue LLMs with substantial knowledge value, necessitating their inclusion under the IP rights of the legal owner. Unfortunately, the service process of LLMs is vulnerable to intellectual property infringement through model extraction attacks, also referred to as model stealing or imitation attacks. Model extraction attacks are akin to knowledge distillation [1], illustrated by the Alpaca case [2]. These attacks create a functionally comparable model in specific domains by distilling the victim model’s knowledge based on the queried victim model’s outputs [3]. Such practices, especially when victim LLMs’ outputs are used for commercial model training, constitute a clear IP infringement. OpenAI’s policy [^1] Terms of use: https://openai.com/policies/terms-of-use, for instance, prohibits using ChatGPT’s outputs for competitive model training, underscoring the legal complexities and the need for stringent IP protection in AI development.
Watermarking technique [4, 5, 6] can embed secret messages into target data, which has been widely used for protecting the intellectual property of the AI models. Existing watermarking methods for IP protection typically follows a similar framework, which utilizes heuristic watermark strategies to embed special signals into the output of the protected model as evidence to identify the suspect model’s IP ownership [7, 8, 9]. For example, He [7] employs lexical modification to embed a watermark in the output text of a targeted model. This watermarked data can then be assimilated through imitation model learning. Zhao [9] implements a unique cosine signal to manipulate the output logits of the model. The logits of every token at each timestep are adjusted to conform to a specific grouped cosine pattern, which the imitation model can then learn. In conclusion, most existing watermark methods for IP protection embed secret message by modifying the distribution of language models through heuristic watermarking rules. However, this inevitably distorts the original distribution of the language model and impacts the normal generation process, resulting in the final generation results deviating from the optimal one and subsequently affecting the model’s performance in its downstream tasks.
Moreover, the attackers can employ diverse adversarial techniques to attack the content generated by the victim model to degrade the effectiveness of the watermark methods. In one hand, attackers are able to modify the watermarked texts, which can reduce the sensitivity of the embedded watermark signals, especially for the watermark methods that rely on fixed synonym replacement rules [7, 8]. In another hand, attackers may also select a part of the watermarked content, and mix them with some clean data to train the imitation model, which can further reduce the impact of the watermarked data on behavior of the imitation model. However, most existing IP protection watermarking methods have not considered the robustness of the designed watermark strategies and the corresponding detection mechanisms, or have only accounted for robustness against specific attacks.
To address these drawbacks, we propose ModelShield, a simple yet effective watermarking method for language model IP protection, which can embed watermarks into model generations without degrading the accuracy of generated content, meanwhile effectively detecting the watermark embedded in the imitation model under the interference of adversarial techniques. Specifically, to avoid distorting the generation distribution of language model via manually-designed rules, we combine the user’s query with a self-watermarking prompt to instruct the victim LLM, which can empower the LLM to adaptively embed watermarks into the generated content. Besides, we further propose a robust watermark detection method to model the strength of watermark signal embedded in content generated by the imitation model and determine the IP infringement. This method also introduces a robust mechanism by filtering the disturbance samples based on the strength of watermark signal encoded in them, to defend against the potential adversarial attacks, like editing or dilution attacks. We conduct extensive experiments on two benchmark QA datasets [10, 11], using ChatGPT as the victim model and GPT-2 large [12], LLama2 [13], and Mistral [14] as the base models for imitation. Results show that our proposed ModelShiled method can adaptively and effectively embed the watermark signals into the generation content, meanwhile reducing the adverse effects of watermarks on the quality of the generated content by 29.08%. Results also demonstrate ModelShied maintains the effectiveness of the watermark while preserving the performance of the language model when adversarial techniques are incorporated to attack the watermarked content.
The contribution of this paper are summarized into three-fold:
  • We introduce a new plug-and-play adaptive watermarking method to combat model extraction attacks without additional training modules. Our method can adaptively embed watermarks messages based on the generated content without manually distorting language model distribution, which can preserve the original performance of language models on downstream tasks.
  • We propose a robust watermark detection algorithm that exhibits high sensitivity to watermark signals embedded within victim models, even when trained on a small subset of user query data (e.g., 400 queries). Our method also maintains its effectiveness despite the presence of various adversarial techniques.
  • Extensive experiments, including scenarios with various watermark detection strategies, different epochs, and robustness tests on two benchmark datasets and three foundational models, demonstrate the effectiveness, harmlessness, and robustness of our ModelShield method.
II Related Works
II-A Model Extraction Attack
Model extraction attacks [15, 16], especially targeting generative models, pose a serious security concern. The attacker aims to create a near-identical replica of the target model by using its outputs. Also known as imitation attacks or model stealing, attackers construct query-response pairs by accessing the victim model’s query API. Subsequently, these pairs are used to train a new model, allowing the imitation to achieve comparable performance to the victim model without needing knowledge of the internal details of the victim model, which renders many backdoor-based watermarking methods ineffective.
II-A1 Types of APIs
Model extraction attacks on language models vary based on the API service type. These attacks target label predictions, output embeddings, or output tokens.
Task-specific label-targeted methods are prevalent [17, 18, 19] at early stage. Attackers exploit the soft labels or hard labels returned by the API to train imitation models. For instance, Krishna [18] developed a model extraction method targeting BERT-APIs for NLP tasks such as sentiment classification and boolean question answering, utilizing the result probabilities as soft labels to train the imitation model. Sanyal [20] proposed a GAN-based framework that trains the imitation model and the generator in tandem to steal the victim model in a hard label setting.
During the era of Embedding as a Service (EaaS), attacks targeting embeddings became prevalent, facilitating a range of natural language processing tasks [16, 21, 22]. Embeddings can provide richer information than labels to assist in training imitation models. Liu [16] initiated the first attack strategy to exfiltrate pre-trained encoders by exploiting the embeddings returned from the victim model. Similarly, Shetty [21] introduced the CSE (Clustering, Selection, Elimination) framework to extract language models from cloud EaaS APIs.
However, the focus has largely shifted towards APIs that directly generate tokens outputs in the era of LLMs [23, 24]. Compared to embeddings, output tokens carry linguistic patterns and contextual dependencies. The imitation model can more easily learn the complex linguistic logic from these information. Wallace [23] launched an attack on a black-box translation system by inputting monolingual sentences and training models to mimic their outputs. Xu [24] demonstrated that black-box APIs, including those that only return tokens, are susceptible to model extraction. The attackers could even outperform the original victim models through unsupervised domain adaptation and multi-victim ensembles. Our research primarily focus on this output tokens-targeted category.
II-B Defense against Model Extraction Attack
To counter model extraction attacks, many efforts have been made to safeguard against model extraction attacks from different perspectives. We categorize defense strategies based on their timing of action into real-time defenses during querying and post-attack IP verification.
II-B1 Real-time defenses
Real-time defenses actively identify behaviors indicative of model theft. This is achieved by analyzing input queries [25, 26] or introducing perturbations to the model’s response [27, 28, 29], thereby complicating the theft process for attackers. PRADA [25] detects abnormal query patterns through statistical analysis of the distance distributions among queries, signaling potential model theft. Additionally, SEAT [26] utilizes a similarity encoder trained with adversarial techniques to spot pairs of similar queries, commonly employed by attackers as they progressively expand a set of seed samples to train the imitation model. Furthermore, defenders can even disseminate misinformation upon detecting out-of-distribution queries [30].
II-B2 post-attack defenses
In contrast, post-attack defenses concentrate on integrating identification markers, such as digital watermarks or fingerprints, into either the victim model or its output. Although post-attack defenses do not prevent the model from being stolen, it plays a pivotal role in asserting IP rights after theft [31, 32]. Watermarking offers a cost-effective and highly efficient approach to counter model extraction attacks. It also can be integrated with real-time defense mechanisms to enhance the overall defense strategy without requiring substantial computational resources. As shown in Fig 12 in supplementary VI-A, we categorize watermarks into four types based on their protection targets: model parameter watermarks, content watermarks, dataset watermarks, and IP protection watermarks. We emphasize that the focus of our work is the IP protection for LLMs against model extraction attack, diverging from parameter watermarking [33, 34, 35, 36] to prevent the theft of the model parameters and content watermarking [37, 38, 39, 40] proving the ownership of model outputs. Specific introductions about the model parameter watermarks, content watermarks and dataset watermarks, as well as the relationships between these watermarks, are presented in the supplementary VI-A.
II-C Language model watermarking method against model extraction attack
The fundamental concept behind language model watermarking against model extraction attacks is to introduce special learnable signals into the outputs of the model to be protected. When attackers use these watermarked model outputs to train their imitation models, these special signals can be detected in the output of the imitation models, allowing for effective IP protection. Based on the watermarking method’s authority to access and manipulate internal computational parameters (such as logits), language model IP protection watermarking methods can be categorized into two distinct classes: black-box watermarking methods and white-box watermarking methods. The detailed workflows of these two types are demonstrated in the supplementary VI-B.
White-box methods require knowledge of the internal details of the model to be protected and the ability to manipulate the text generation process. The representative white-box method is GINSEW [9], which groups and manipulates the conditional token probabilities of language model. It embeds specific frequency sine perturbations as watermark signals in the grouped probabilities, enabling the imitation models to learn this signal pattern during training. Watermark detection involves querying the imitation model for token probabilities, conducting matching filtering to detect grouped conditional probability outputs, and checking for the presence of embedded specific frequency signals. Similarly, the EmbMarker [41] uses a set of medium-frequency trigger words as watermark signals. EmbMarker selects a target embedding as the watermark and inserts it into the embeddings of texts containing trigger words as a backdoor. However, it is not suitable for the current LLM market compared to the previous era of EaaS.
Black-box methods, on the other hand, do not require knowledge of the internal details of the model to be protected and can provide more universal protection for language models. A classic work in the black-box domain is the lexical watermarking proposed by He [7] for protecting intellectual property. This method involves synonym replacement and English-American grammar substitution in the output of the language model to be protected. Specifically, a set of words $W$ is selected, and for each word $w$ in $W$, a semantically equivalent substitution is identified and used to replace $w$ in the output of the victim model. The objective is to alter the word distribution such that the imitation model learns this skewed pattern. Attackers, when training their own imitation models using the replaced data, tend to output these substituted words compared to a legitimate model. Taking it a step further, the CATER [8] refines the synonym replacement strategy, aiming to reduce overall word distribution distortion while enhancing the change in conditional word selections. However, it requires the conditional probability of the victim language model and is no longer suitable for black-box scenarios. Similarly, Li [42] replaces tokens with their synonyms after language model generation, protecting the IP of language models in code generation scenarios. Our watermarking method also belongs to black-box methods.
III Methodology
Figure 2: The workflow of our watermarking method. Our watermarking method involves two main processes: embedding and extraction, indicated by blue and yellow frames, respectively. In the embedding phase, a user’s query is combined with an automatic watermark generation instruction as input to the victim model, which then produces watermarked text outputs. These outputs can be used by malicious users to train imitation models; To verify if suspect models were trained using the victim model’s outputs, we analyze the query history between the suspect model owner and the victim model and conduct watermark detection on the suspect model’s text outputs. If a t-test on the top 1% of data with high watermark scores results in a p-value below 0.05, we conclude that the suspect model is an imitation trained with watermarked data.
III-A Preliminary
Terms commonly used in language model extraction attack scenarios include:
  1. 1.
    Victim Model $M_{V}$: The target of the model extraction attack, where the attacker can obtain the output from the victim’s model; it is also the object protected by our watermarking method, also known as the protected model.
  2. 2.
    Imitation Model $M_{I}$: An imitation model is a model trained with data from a victim model with the goal of becoming a replica of the victim model that is comparable in performance to the victim model in some way, which is also referred to as a thief model, mimic model.
  3. 3.
    Backbone Model $M_{O}$: Given the high costs associated with training large language models from scratch, model extraction attackers generally utilize data from victim models to perform fine-tuning on existing open-source base models to create imitation models. We refer to the base models used by these imitation models as foundation models, base models or backbone models.
  4. 4.
    Legitimate Model $M_{L}$: It refers to one that is trained on non-watermarked data using a foundational model. Although it may achieve performance similar to the victim model, it is distinct in that it did not utilize the outputs of the victim model for training.
  5. 5.
    Suspect Model $M_{S}$: The suspected model is the subject of watermark detection. It might be an imitation model or a legitimate model. The watermark detection process aims to discern whether the suspected model has used watermarked data from a victim model for training.
  6. 6.
    Watermarked Output $W_{W}$: Watermarked outputs refer to the texts generated by a protected model after incorporating a watermarking mechanism, also known as watermarked data. This data, which attackers can obtain, may be used to train imitation models. Attackers have the ability to edit and modify this watermarked data.
  7. 7.
    Suspected Output $W_{S}$: The outputs of a suspected model are the basis for watermark detection decisions. If the suspected model is an imitation model, its outputs will exhibit a preference for the watermarks.
III-B The basic property of language model IP protection watermarks
We propose four fundamental properties crucial for effective watermarks in model extraction attack scenarios:
1. Effectiveness: Given the outputs from a suspect model, the owner of the victim model can determine within a limited time whether the suspect model is an imitation of the victim model.
2. Learnability: The watermark should be integrable through imitation model training, ensuring its effectiveness and detectability even after model training.
3. Harmlessness: The watermark should have minimal impact on the original language model’s performance, thereby preserving the user’s normal query experience while minimizing the attacker’s awareness of its presence.
4. Robustness: Watermarks should possess a certain degree of resilience to withstand attackers’ edits and modifications that maintain semantic consistency in watermarked data. Additionally, they should withstand attempts by attackers to dilute watermarks presence during the training of imitation models.
III-C Adaptive Self-Watermarking Mechanism
Our goal is to devise a straightforward yet effective method for automatic watermark generation, avoiding manual intervention and extensive training processes like pre-training or fine-tuning. Given that existing LLMs possess an almost human-like capacity for understanding and adhering to instructions [43, 44], our strategy leverages this strong instruction-following ability as a defensive mechanism in system mode. By utilizing the inherent proficiency of LLMs in token selection [45], we aim to guide them autonomously in generating watermarks to defend model extraction attacks. The watermark generation process resembles a person reminding themselves to incorporate a unique watermark known only to them during the spontaneous creation stage [46, 47].
Building on this insight, we introduce an additive system-mode self-watermarking method. We explore the capability of LLMs to autonomously and independently generate watermarks. Our method utilizes a system-generated prompt that instructs the LLM to embed an appropriate watermark based on its own assessment. For example, we implement a system prompt that integrates the user query in the following manner:
[USER QUERY]: Write a email to your boss explaining why you should have a pay raise. [SYSTEM PROMPT]: Use specific strategies and include some special words as watermarks in your generated text. Ensure these watermarks are present in your answer.
This system-mode self-watermarking approach is readily applicable across various LLMs, serving as an efficient identifier against model extraction without compromising the original output quality of the LLMs. In our experiments with several similar prompts, we observed comparable watermark generation effects as shown in the Table I. We find that due to the language model’s powerful prompt understanding capabilities, our automatic watermarking method is insensitive to the specific prompt used. While optimizing the automatic prompts might enhance performance in other downstream tasks, we don’t consider prompt searching as the focus of our work. Clearly, articulated prompts with similar intentions can also achieve similar effects in generating watermark words. Afterward, the model owner can format the generated text, returning only relevant answers to the user and recording the watermark words from the query.
Prompt 1 Insert watermark words only you know into your responses. Here’s the question: ### {User Query}
Prompt 2 Embed some special watermark words in your generated texts. Ensure their presence in your response to my upcoming question. Here’s the question: ### {User Query}
Prompt 3 Embed watermarks in your text responses. Here’s the question: ### {User Query}
Prompt 4 Use some special words as watermarks in your generated text and tell me in the end. Here is my question: ### {User Query}
Prompt 5 Incorporate specific watermark words in your text and notify me when your generation is done. Now, here’s my question: ### {User Query}
Prompt 6 Insert specific watermark words in your text and notify me at the end. My question is: ### {User Query}
TABLE I: System instructions for enabling the victim model to generate watermarks automatically.
III-D Robust IP Infringement Detection
When model owners suspect a model theft, firstly they can use their prior knowledge to detect whether the theft queries their LLM services before. It is a common assumption in model stealing verification followed by [9] , also there is evidence that OpenAI [48] actually has the query history of every user’s query and answers that OpenAI suspended ByteDance’s account after it used ChatGPT to train its own competing AI model.
When an individual is suspected of using query data to develop their own competing imitation model, the model owner can retrieve the suspect’s query records for further detection. After that, the model owner can use the suspected theft’s history query-response $\{Q,A\}$ to detect the suspicious model $M_{S}$. If the theft aims to build a functionally similar model using the watermarked data $A$, the imitation model $M_{I}$ will learn from this data and generate more watermarked tokens than the innocent foundation model $M_{O}$ or the foundation model trained with normal data $M_{L}$. In real-world scenarios, an attacker might not utilize the entire watermarked dataset $A$ from a specific victim model to train their imitation model. Let’s denote the dataset used to train the imitation model as $D$. The attacker may only use a portion of $A$ for fine-tuning their imitation model, which we denote as $A^{D}$. It is evident that $A^{D}$ is a subset of $D$ (expressed as $A^{D}\subseteq D$), and $A^{D}$ is also a subset of $A$ (expressed as $A^{D}\subseteq A$). We define the ratio $\rho$ and the ratio $\eta$ as:
$\rho=\frac{|A^{D}|}{|D|},\ \eta=\frac{|A^{D}|}{|A|}.$ (1)
For the attacker, both ratios $\rho$ and $\eta$, are known and controllable. However, the watermark detector has no knowledge of these two ratios.
Formally, the attacker has access to $\{Q,A,D,Q^{D}\}$, while the watermark detector is informed of $\{Q,A,WM,Y_{S}\}$ ,where $WM$ represents the watermark sets recorded by the model owner and $Y_{S}$ denotes the outputs from the imitation model using the query sets $Q$.
Detecting IP violations in model extraction attacks requires methods to identify faint yet statistically significant watermark signals amid a large volume of tokens. We employ the hypothesis testing method to decide if the suspect model $M_{S}$ has used the watermarked data for training. The null hypothesis, $\mathcal{H}_{0}$, posits that the suspect model $M_{S}$ is not trained using the watermarked data sourced from the victim model $M_{V}$. Consequently, it implies that $M_{S}$ should exhibit no particular affinity for watermarks, resulting in watermark scores for the texts output by $M_{S}$ not being significantly elevated. we have $1-\alpha$ confidence to claim that $M_{S}$ is a replica of victim model $M_{V}$ if the statistical test is able to reject $\mathcal{H}_{0}$ at a significance level (p-value) smaller than $\alpha$ (normally we take $\alpha=0.05$). In other words, If we reject $\mathcal{H}_{0}$ with a smaller p-value, it means we detect watermarks and we have higher confidence that the suspect model did use $A$ from the victim model $M_{V}$ for training.
In practical black-box testing scenarios, where we only access the suspect model’s outputs, we use a fast, prior-based watermark verification method. Should we gain access to the base information of the imitation model, we proceed with a more detailed contrastive verification.
Prior Rapid Watermark Verification:
This method rapidly judges based on the outputs $W_{S}$ of the suspect model $M_{S}$ to $Q$. It enables quick and effective watermark identification without relying on machine learning modules.
Detailed Contrastive Watermark Verification:
With knowledge of the suspect model’s base model information, we can further compare the output differences between the suspect model and the base model using the Kolmogorov-Smirnov (KS) test [49]. Furthermore, watermark detectors can train the base model using unwatermarked data on the known base model, allowing for a more detailed comparison with the suspect model.
In prior rapid verification, we only need the outputs $Y_{s}=\{y_{0},...y_{i},...y_{n}\}$ of the $M_{s}$ responded to $Q=\{q_{0},...q_{i},...q_{n}\}$ to prove that the word distribution in texts generated by $M_{I}$ is skewed towards $M_{V}$’s output $W_{W}$.
Firstly, we define the Sentence Watermark Score (SWS) $S_{i}$ for each text $y_{i}$ generated by the suspect model:
$S_{i}=\frac{\sum_{w\in WM_{i}}\mathbbm{1}_{{y_{i}}}(w)\times{l}(w)}{l(y_{i})}$ (2)
where $WM_{i}$ is the watermark word set for query $q_{i}$ and $\mathbbm{1}_{y_{i}}(w)$ is the indicator function for the presence of $w$ in $y_{i}$. $l(\cdot)$ is the function to count tokens length. In other words, we calculate the total length of all unique watermark words found in the suspect model’s generated text, normalized by the length of the generated text. The Average Sentence Watermark Score (ASWS) is calculated by ${\sum_{i=1}^{n}{S_{i}}}/{n}$.
The Watermark Score (WS) for a collection of texts $Y={\{y_{0},...y_{i},...y_{n}\}}$ is:
$WS(Y)=\frac{1}{k}\sum_{i=0}^{k}S_{\sigma(i)}(y_{i}),$ (3)
where $\sigma(\cdot)$ is the descending order function, and $S_{\sigma(i)}$ is the Sentence Watermark Score of generated text $y_{i}$. we compute the top $k$ highest Sentence Watermark Score sentences of $Y$, $k=max{(0.01\times|Y|,100)}$. We dynamically determine the value of $k$ based on the number of samples $n$ we can obtain and the desired sensitivity level.
We define a threshold $\theta$ to assess whether the suspect model $M_{s}$ imitates the victim model $M_{V}$. This threshold is derived from the precomputed Watermark Score $WS(Y_{h})$ of human texts without watermarks. Our method’s adaptability, distinct from [7]’s approach, precludes knowing watermark word ranges beforehand. Therefore, we initially calculate the Watermark Score for a collection of general domain human texts $Y_{h}$ representing a range of typical, non-watermarked human discourse. The threshold $\theta$ is set by averaging these Watermark Score and adding a relaxation coefficient $\gamma$ to surpass the highest watermark score in each domain, minimizing accidental watermark activation in normal texts. Thus, $\theta$ accounts for both average tendencies and score variances. The threshold is defined as:
$\theta=\frac{1}{m}\sum_{j=0}^{m}{WS(Y_{h}^{j})}+\gamma$ (4)
The detailed calculation of human dataset watermark scores and threshold determination are discussed in IV-A. We then apply a one-sided one-sample t-test [50] to evaluate p-values from this test against $\theta$, conducted on the top $k$ sentences from the suspect model $M_{S}$, ranked by Sentence Watermark Scores. A p-value below $\alpha$ (usually 0.05) allows us to reject the null hypothesis $H_{0}$ with confidence.
(a) Using Gpt2-Large as the base model on HC3
(b) Using Llama2 as the base model on HC3
(c) Using Mistral as the base model on HC3
(d) Using Gpt2-Large as the base model on WILD
(e) Using Llama2 as the base model on WILD
(f) Using Mistral as the base model on WILD
Figure 3: Main Experiment Results: We trained imitation models ($M_{I}$) using 4000 watermarked data based on three different foundational models, repeating the process ten times. The results demonstrate the average sentence watermark score from all texts generated by the imitation models, as well as their question-answering performance. Compared to outputs from the original foundational models ($M_{O}$) and models trained with normal data ($M_{L}$), the watermark scores of the imitation models were significantly higher, while their question-answering performance remained comparable to that of normal models trained without watermarked data ($M_{L}$).
In the detailed contrastive watermark verification, we employ two-sample KS test [49], which is inspired by the detection method in Membership Inference Attacks [51]. The KS statistic is the maximum difference between the cumulative distribution functions (CDFs) of the two samples. Given the distribution $D_{S}$ of Sentence Watermark Scores from the outputs of suspect model $M_{S}$, the distribution $D_{L}$ from the model trained with unwatermarked data $M_{L}$, the distribution $D_{O}$ from the origin base language model $M_{O}$, we intuitively think the Sentence Watermark Score distribution of the imitation model $M_{I}$ should have apparent difference from $M_{L}$ or $M_{O}$. The KS test evaluates whether these distributions exhibit significant differences, with an emphasis on their cumulative properties. A low p-value from the KS test indicates that the distributions are significantly different, while a high p-value indicates that the distributions are similar and consistent with each other. the KS statistic defined for $M_{L}$ ,$KS_{L}$ is defined as:
$KS_{L}=\sup_{x}|F_{D_{S}}(x)-F_{D_{L}}(x)|$ (5)
the KS statistic defined for $M_{O}$, $KS_{O}$ is defined as:
$KS_{O}=\sup_{x}|F_{D_{S}}(x)-F_{D_{O}}(x)|$ (6)
where $sup$ denotes the supremum (the least upper bound) and $F_{D_{S}}(x)$, $F_{D_{L}}(x)$ and $F_{D_{O}}(x)$ are the CDFs of the distribution $D_{S}$ ,$D_{L}$ and $D_{O}$, respectively.
If the $KS_{{L}}$ and $KS_{{O}}$ are both high, which set the p-value of the test, we can deduce that the watermark score distribution of the suspect model $M_{S}$ markedly differs from those of the non-watermarked data finetuned model $M_{L}$ and the original model $M_{O}$, which can further conclude that the suspect model $M_{S}$ is an imitation model of victim model $M_{V}$.
IV Experiments and Analysis
(a) Gpt2-hc3
(b) Llama2-hc3
(c) Mistral-hc3
(d) Gpt2-wild
(e) Llama2-wild
(f) Mistral-wild
Figure 4: Cumulative Distribution Function (CDF) of Sentence Watermark Scores: This graph shows the CDF of 4000 sentence watermark scores of 4000 samples for $M_{I}$, $M_{L}$, and $M_{O}$. It clearly illustrates that the distribution of the Imitation model is significantly distinct from the other two types.
IV-A Self-Generated Watermark Datasets
As we are the first to explore the self-generated watermarking capability of language models, we build two LLM self-generated watermark datasets. Given the widespread utilization of question-and-answer interactions in language model services, we built the self-generated watermarks dataset based on two publicly available open-ended question-answering datasets, namely Human ChatGPT Comparison Corpus (HC3) [10] and InstructWild (WILD) [11]. HC3 comprises comparison responses from both human experts and ChatGPT across diverse domains, including finance, medicine, law, and psychology. WILD, on the other hand, is an open QA mind strom dataset rich in varied and imaginative instructions. To simulate the process of a model knowledge thief querying victim language models, we sampled 4,000 instances from each of these datasets and employed our method to embed watermarks into the answers provided by the victim models [^2] Our data and models are available at https://github.com/amaoku/ModelShield . Then, based on the constructed watermark dataset, we set the watermark threshold for the upcoming experiments. As shown in the Table II, we calculate Watermark Scores for human texts in each domain based on the union of watermark words generated on watermarked datasets. This includes analyses of human responses within HC3 and WILD, where they closely match the domain and are most likely to trigger watermark words unintentionally. For each dataset, we selected texts with the top 1% Sentence Watermark Scores to compute their Watermark Score. In our experiments, we usually set threshold $\theta$ at 0.11.
Human Dataset Numbers Watermark Score
TWEET [52] 2588579 $0.0692_{0.0279}$
NEWS [53] 1713999 $0.0373_{0.0137}$
MOVIE [54] 1039403 $0.0654_{0.0359}$
FINANCE [55] 68912 $0.0001_{0.0010}$
GUNANACO [56] 53461 $0.0001_{0.0019}$
SENTIMENT [57] 50000 $0.0101_{0.0232}$
HC3 [10] 37175 $0.0993_{0.0279}$
WILD [11] 52191 $0.1005_{0.0849}$
TABLE II: Watermark Score on human dataset. We aggregate the sentence watermark scores of the top 1% of sentences in the dataset to statistically analyze the watermark scores.
IV-B Model Extraction Attack Simulation
We simulate the process of stealing knowledge from a victim model by fine-tuning its output using a foundational model. We follow the strictest setting that we access only the output tokens of the black-box language model.
In our study, we employ ChatGPT [58] as the victim model, querying it and applying our method to generate answers containing watermark words. We use the transformer-based [59] model as the imitation model due to its widespread adoption and effectiveness in text generation. We utilized GPT-2 Large-0.7B [12], Llama2-7B [13], and Mistral-7B [14] as backbones for our imitation models. These selections represent a diverse set of models: GPT-2 Large as a homologous model and LLaMA-7B and Mistral-7B as heterogeneous large language models.
To simulate the process of a model extraction attack, we utilize the watermarked data generated by GPT-3.5 to fine-tune the backbone model, subsequently employing the fine-tuned model as an imitation model. Additionally, we used the same query with GPT-3.5, but without any additional watermark requirements, to generate data without watermarks. We then used these watermark-free data to train the backbone model $M_{L}$ as a contrast. The outputs from the model $M_{L}$ trained with the watermark-free data and the original backbone model $M_{O}$’s outputs are used as a comparison for the outputs of the imitation model $M_{I}$ trained with watermarked data.
Details of Imitation Model Training: The imitation model training details are listed in Table III. The experiments with Gpt2-Large are conducted on NVIDIA 3090, while those with Llama2 and Mistral are conducted on NVIDIA A800. We conducted full-parameter fine-tuning on GPT-2 and Llama2 without using lightweight methods, while Mistral undergoes fine-tuning using the lightweight LoRA [60]. Our findings indicate that our watermarks could be effectively learned under both scenarios.
Base model Gpt2-Large Llama2 Mistral + LoRA
Batch size 1 1 1
Max learning rate 1e-5 2e-5 2e-5
Traninable params 774,030,080 6,738,415,616 41,943,040
Cutoff length 1024 1024 1024
Optimizer Adam Adam Adam
Epochs in epoch test in epoch test 10
Warmup steps 10 10 10
LoRA rank - - 16
LoRA alpha - - 32
LoRA dropout - - 0.05
TABLE III: Training hyperparameters of imitation model.
IV-C Metrics
The effectiveness and learnability are gauged using the sentence watermark score or watermark score, as elucidated in our watermark verification method. Robustness is evaluated by whether the watermark can retain its effectiveness after being attacked. Harmlessness is evaluated by the impact of the watermark on the language model’s performance in its intended tasks. We evaluated the harmlessness from two perspectives: firstly, the harmlessness of the watermarked text, where we observed if there was any decrease in output linguistic quality from the victim model ($M_{V}$) after incorporating the watermark mechanism. Secondly, we evaluate harmlessness by comparing the task performance of the imitation model ($M_{I}$) trained with watermarked data against a model ($M_{L}$) trained with non-watermarked data. Specifically, we assess the language model’s QA task performance using BLEU [61] and ROUGE [62] metrics based on human gold standards. The specific calculation formulas of BLEU and ROUGE we used is in supplementary VI-C.
IV-D Performance Evaluation
As shown in Fig 3, we evaluated the effectiveness and harmlessness of watermarks on models fine-tuned with watermarked data ($M_{I}$), with non-watermarked data ($M_{L}$), and original base models ($M_{O}$) using 4000 samples on two datasets and three foundational bases. The left bar in our representation indicates the effectiveness of the watermark, quantified specifically by the average Sentence Watermark Score of all texts generated by $M_{S}$. The right bars illustrate the non-intrusiveness of watermarks, reflected through task performance metrics.
Firstly, it is evident from the data that the watermark score of the $M_{I}$ is statistically higher than that of the $M_{L}$ and the $M_{O}$. This finding suggests a distinctive difference in $M_{I}$, which has been trained on watermarked data. We repeated 10 times experiments and conducted further detailed contrastive verification KS tests, as Table IV illustrates.
We consistently observed that the p-values of statistical tests between $M_{I}$ and both $M_{O}$ and $M_{L}$ are significantly less than $10^{-12}$. In 10 repeated experiments, each with 4000 samples, we did not observe instances where the p-value for the imitation model ($M_{I}$) exceeded 0.05 (false negative) or where the p-values for the backbone model ($M_{O}$) and the legitimate model ($M_{L}$) fell below 0.05 (false positive). Furthermore, rapid watermark verification methods yielded p-values consistently below $10^{-9}$, confirming our watermark method’s reliability.
In Fig 4, we plotted the cumulative probability distribution of watermark scores for the three types of models, further reflecting the significant statistical differences between $M_{I}$ and both $M_{L}$ and $M_{O}$. This result strongly suggests that $M_{I}$ has effectively learned the watermark patterns, showing a marked preference for watermark tokens.
[…]
HC3 [10] Gpt2-Large 1.73e-15 2.13e-14 2.31e-13
Llama2 9.49e-64 3.27e-14 2.41e-29
Mistral 6.25e-10 7.12e-179 2.29e-110
WILD [11] Gpt2-Large 6.23e-68 4.64e-12 7.04e-17
Llama2 1.20e-57 5.72e-13 9.85e-43
Mistral 5.64e-18 5.53e-214 3.17e-122
TABLE IV: Average P-values of Watermark Detection in Rapid and Detailed Contrastive Verification. In Rapid Verification, only the outputs from the suspect model are used. In Detailed Contrastive Verification, the detector has information about the backbone model ($M_{O}$) used for the model extraction attack. This allows for a comparison with the original backbone model $M_{O}$ or legitimate model $M_{L}$.
[…]
HC3 [10] w/ 0.149 0.079 9.94
w/o 0.146 0.075 9.86
WILD [11] w/ 0.356 0.194 13.37
w/o 0.350 0.206 12.78
TABLE V: Task performance of the victim modelon the HC3 and WILD datasets for question-answering tasks before (w/o) and after (w/) the addition of our watermarking mechanism.
As Table V depicts, for the victim model $M_{V}$, the implementation of our watermarking mechanism shows negligible differences in performance on QA tasks, indicating minimal impact from system-mode self-watermarking prompts. This is attributed to the fact that our instruction-guided method avoids forcefully distorting the original distribution of the language model. As for the imitation model, we observe that the quality of answers produced by the $M_{I}$ is on the same scale as that of the $M_{L}$, and at times, it even surpasses the answer quality of the $M_{L}$, signifying that the training with watermarked data doesn’t impede the model’s capabilities. This subtlety ensures that the attackers do not easily detect anomalies or even engage in reverse engineering attacks against the watermarks.
IV-E Effectiveness of Different Watermark Detection Strategies
(a) Different watermark strategy for Gpt2
(b) Different watermark strategy for Llama2
(c) Different watermark strategy for Mistral
Figure 5: Different watermark strategy. We tested the average sentence watermark score of three base models (Gpt2, Llama2, and Mistral) on the HC3 and WILD datasets.
We adopt three distinct criteria to examine the watermarks produced by the model from various perspectives.
Firstly, we consider all watermark words indiscriminately generated by the model, designated as $WM$. This broad approach allows us to assess the model’s overall capability in generating watermarks.
Secondly, we focus on watermark words that are not present in the query, referred to as $WM_{Q}$. This criterion is pivotal as it aligns with our goal of enabling the language model to generate watermarks naturally and autonomously, without altering its original output distribution. Therefore, it’s essential that the watermark generation is not influenced severely by the query’s content. By definition, $WM_{Q}$ forms a subset of the total watermarks $WM$.
The third criterion identifies high-entropy watermarks that are not present in the query and exhibit low frequency or rarity. We posit that such high-entropy words, usually more complex for both humans and language models, are particularly suited for watermarking purposes. This category naturally forms a subset of $WM_{Q}$.
Specifically, we tokenize HC3 and WILD human answers and count the frequency of each token. Watermarks that are absent from the query and appear less than five times within the human response corpus are classified as $WM_{H}$. The relationship among these watermark categories is described by the equation:
$\{WM_{H}\}\subseteq\{WM_{Q}\}\subseteq\{WM\}$ (7)
Generally, the size of $WM_{H}$ is the smallest when compared to $WM_{Q}$ or $WM$. Consequently, the ASWS for $WM_{H}$ tends to be relatively lower as shown in the Fig 5.
IV-F Performance Under Varying Training Adequacies
(a) Epoch Testing with Gpt2-Large as the Base Model on HC3
(b) Epoch Testing with Gpt2-Large as the Base Model on WILD
(c) Epoch Testing with Llama2 as the Base Model on HC3
(d) Epoch Testing with Llama2 as the Base Model on WILD
Figure 6: Epochs Analysis: We examined the efficacy of watermark detection in imitation models trained with watermarked data across different training epochs, utilizing three watermark strategies.
We evaluated the effect of different training epochs on watermark detection in imitation models, using both Gpt2 and Llama2, as illustrated in Fig 6. Firstly, it is clear that as the number of training epochs for the imitation model increases, the Average Sentence Watermark Score in detection consistently rises. This indicates that if an attacker aims to replicate the performance of the victim model on downstream tasks by training the imitation model over multiple epochs, the watermark will be steadily and consistently learned.
The speed of performance convergence varies among different types of models. For models with a smaller number of parameters, such as Gpt2, both the watermark detection rate and model performance can steadily increase with more training epochs. Notably, the rate of watermark convergence is faster than the convergence rate of the model’s performance on downstream question-answering tasks. However, for models like Llama2, which are pre-trained on massive datasets, excessive fine-tuning epochs do not further enhance the model’s performance on its original tasks, but watermarks can still be detected reliably and effectively.
IV-G Comparison with baseline
Our comparative method primarily revolves around watermarking methods for language model IP protection. Given our aim to minimally impact the language model’s performance, we have restricted access to token probability manipulation and only receive the output token of the target model (ChatGPT). Consequently, our comparison is constrained to the work conducted by He [7] in this particular setting. The baseline method involves embedding watermarks in the output content of the model to be protected through post-processing. Specifically, this method mainly includes synonym replacement in the output content of the model to be protected. Following this, attackers utilize the altered data to train their own models. When compared to models trained on normal data without knowledge of synonym modification rules, models trained with watermarked data tend to exhibit a preference for synonymous with the embedded watermark.
Method Generated sentence
Original model Traditionally, new seasons of TV shows have premier- ed in the fall because that is when the major networks have their ”upfronts”…
Baseline [7] Traditionally, novel seasons of TV shows have premier- ed in the fall because that is when the major networks have their ”upfronts”…
Ours The reason TV shows are usually in the fall is because it’s when the TV networks release their new and excit- ing shows, which creates a lot of nuzz and excit- ement amomg viewers.
TABLE VI: Examples of texts provided for users comparing with baseline on HC3. Synonyms as watermarked word will cause weird generation in some case.
(a) comp. w. baseline [7] on HC3
(b) comp. w. baseline [7] on WILD
Figure 7: Compare with baseline: We compared the linguistic quality (indicated by PPL) and task performance (indicated by ROUGE and BLEU scores) of texts provided for users after our method and baseline methods embedded watermarks in the output of the victim model. The baseline method’s rigid synonym substitution resulted in a loss of quality in the watermarked text outputs.
(a) Performance Evaluation of Using Gpt2-Large on HC3
(b) Performance Evaluation of Using Llama2 on HC3
(c) Performance Evaluation of Using Gpt2-Large on WILD
(d) Performance Evaluation of Using Llama2 on WILD
(e) Detailed Distribution of SWS using Gpt2-Large on HC3
(f) Detailed Distribution of SWS using Llama2 on HC3
(g) Detailed Distribution of SWS using Gpt2-Large on WILD
(h) Detailed Distribution of SWS using Llama2 on WILD
Figure 8: Robustness Test against Editing Attack. In figures (a)-(d), we present the watermark detection efficiency (represented by the average sentence watermark score of all imitation model outputs) and task performance (measured by ROUGE) across various imitation models and watermarked data post-editing attacks. Figures (e)-(h) display a more detailed distribution of sentence watermark scores, where we plotted the distribution for the top 1% highest scoring samples. The results show that even after editing attacks, the watermark scores in the imitation model outputs remain significantly distinct from those of normal models trained without watermarked data and the base models.
We have compared ModelShield with the baseline in three aspects: watermark detection, the linguistic quality of the protected model’s output, and performance in downstream tasks. As for watermark detection, the baseline method yields a p-value of less than $10^{-8}$ using the default setting, while ModelShield achieves a p-value of less than $10^{-17}$. Both results are below the $0.05$ threshold, satisfying the criteria for rejecting the null hypothesis.
Due to the different watermark detecting algorithms, we evaluate the output’s linguistic quality and performance in downstream tasks of the protected model by contrasting the baseline method at similar p-value magnitudes (around $10^{-8}$). As the Fig 7 shows, our method can provide users with a better linguistic quality response (15.32% and 13.80% PPL $\downarrow$) and a higher QA score in downstream tasks (27.27% and 29.08% ROUGE $\uparrow$). In certain situations, the baseline method can significantly reduce the original output quality of the model, as illustrated in the Table VI. Forcibly replacing synonyms leads to unsuitable expressions in some cases.
(a) SWS Distribution for GPT-2 Large on HC3
(b) SWS Distribution for GPT-2 Large on WILD
(c) SWS Distribution for Llama2 Large on HC3
(d) SWS Distribution for Llama2 Large on WILD
Figure 9: Attackers select watermarked data subsets for imitation model training. The horizontal axis represents the ratio $\eta$ of watermark data used for training the imitation model out of the total watermark data.
IV-H Robustness to Adversarial Attacks
We simulate three adversarial attack that pose challenges to watermark detection in realistic scenarios: user editing attacks, selective training on watermarked data subsets, and training with a mixture of watermarked and unwatermarked data.
IV-H1 Editing Attack
We tested the resilience of our watermarking method against attackers who modify watermarked texts. Specifically, we conducted experiments in four scenarios: random insertion, random deletion, random replacement, and rearranging token order within sentences while maintaining semantic consistency (random scramble). Details of these attacks are given below:
1. Random Insertion: This involves inserting a random token at a random location within the same phrase. 2. Random Deletion: A random token is selected from the phrase and completely removed. 3. Random Replacement: In this attack, a random token is selected and substituted with another randomly picked token. 4. Random Scramble: Two tokens in the sentence are chosen at random and swapped, altering the sequence but not the overall meaning.
We can initially confirm that attacking watermarked data will undoubtedly impact the performance of watermark detection. However, as observed from the Fig 8, even after various editing operations on watermarked data, the imitation models trained using our watermarking method still show a statistically significant difference in sentence watermark score compared to $M_{L}$ and $M_{O}$. Moreover, our watermarking method does not compromise the original task performance of the model despite being subjected to editing attacks. The task performance of the imitation model remains comparably high, akin to scenarios without such attacks.
IV-H2 Selective Training on Watermarked Data Subsets
(a) SWS Distribution for Llama2 on HC3
(b) SWS Distribution for Llama2 on WILD
Figure 10: Attackers combine non-watermarked data from other sources for training imitation models: In this scenario, attackers use all the watermark data obtained from queries and mix in non-watermarked data from other sources to train imitation models. We controlled the watermark data at a fixed value (400 entries) and mixed in non-watermarked data, varying the watermark data’s proportion $\rho$ to 10%, 25%, 50% and 75%.
In this mixed training scenario, we prepared six distinct datasets, each containing a varying ratio $\eta$. The values of $\eta$ were set at 0, 0.1, 0.25, 0.5, and 0.75, and we ensured that the size of the total watermark dataset $|A|$, consistently remained at 4000 instances. The outputs of the suspect model were then evaluated based on our comprehensive knowledge of the query history.
(a) BertScore and Sentence Watermark Score Using Gpt2 on HC3
(b) BertScore and Sentence Watermark Score Using Gpt2 on WILD
(c) BertScore and Sentence Watermark Score Using Llama2 on HC3
(d) BertScore and Sentence Watermark Score Using Llama2 on WILD
Figure 11: Scatter Plot of Watermark Scores and BertScore: Yellow dots represent outputs from imitation models, while blue dots indicate outputs from the original base models, with the intensity of the color indicating the sentence watermark score. Each sample is evaluated for its similarity to the watermarked output of the victim model (under the same query) using BertScore. The x-axis represents BertScore’s Precision, and the y-axis shows BertScore’s Recall. The closer the points move towards the upper-right corner, the more similar the text is to the watermarked output of the victim model.
Even though attackers use only a portion of the watermarked data for training, we conduct watermark detection on all the queries $Q$ known to have been scraped by the victim model’s owner. As depicted in Fig 9, a distinct statistical discrepancy is apparent between models trained with and without watermarked data (with a p-value of about $10^{-2}$), even when only 10% of the watermark data (400 samples) are used for imitation model training.
This not only demonstrates the robustness of our watermark verification method but also indicates the learnable nature of the watermarked data. It can also be seen that as $\eta$ increases, the average sentence watermark score rises accordingly. This aligns with our intuition that the more watermarked data used for training the imitation model, the easier it is to perform watermark detection.
IV-H3 Mixture of unwatermarked data
In a scenario similar to selective training on watermarked data subsets, attackers might mix all queried data from the victim model with additional unwatermarked source data for training the imitation model. We set $A^{D}=A$ in this case. Our experiments maintain a fixed number of watermarked data while varying the total training data $|D|$, altering $\rho$ to values of 0.1, 0.25, 0.5 and 0.75. We only detect watermarks on queries known to the victim model’s owner since the attacker can blend in any unwatermarked data. The results, as illustrated in Fig 10, reveal that increasing the proportion of watermarked data correspondingly elevates the watermark detection in the text generated by the imitation model. Watermarks are detectable in all cases with p-values below $10^{-2}$, without affecting the performance on downstream QA tasks.
IV-I LLM Generation Visualization
We utilize watermarks to verify the IP copyright of language models, primarily to ascertain if a model’s output has been used for unauthorized training. Through the following experiments, we will show the alignment between the watermarked data and the victim model’s output, indicating that detecting watermarks is a reliable indicator of the victim model’s output being used in training.
As illustrated in Fig 11, we analyze the BertScore [63] comparisons between the imitation model’s outputs and the original foundational model’s responses to identical queries. It’s evident that the output of the imitation model, in various text relevance scenarios, consistently shows high sentence watermark scores, proving that the watermark can effectively transfer to the imitation model and be effectively detected, making it a necessary mechanism. Conversely, outputs from base models (represented in blue) show generally low sentence watermark scores, proving that the watermark is not prone to false positives.
Moreover, we noted that the data points with elevated watermark scores (illustrated in dark yellow) are evenly distributed throughout the imitation model’s outputs. This suggests that the watermarks generated by the language model are not anomalous deviations from typical text but rather share a notable resemblance to normal texts. This aspect of the watermarks lends them a certain level of non-distinctiveness.
Furthermore, the output of the imitation model does not align perfectly with that of the victim model, as high watermark scores (dark yellow) do not consistently cluster in the top-right corner of the graph. Many samples exhibit partial overlap, illustrating that direct matching based on the victim model’s outputs introduces significant noise.
Additionally, the outputs of the imitation model generally resemble those of the victim model, corroborating the idea that the inherent learnability of watermarks serves as an effective indicator for detecting model extraction attacks.
V Conclusion
In this paper, we propose ModelShield, a plug-and-play watermarking method for language models designed to combat model extraction attacks. Unlike existing IP protection watermarking methods that modify model’s output or distribution with heuristic rules, ModelShield uses system prompt guidance to minimize distortion seamlessly. Leveraging the self-reminding capabilities of LLMs, ModelShield integrates user queries with a self-generated watermark system prompt, subtly embedding watermark words into outputs without altering the model’s internal probabilities or outputs. This approach maintains text quality while ensuring watermark effectiveness and robustness. Extensive experiments validate ModelShield’s efficacy, showing remarkable detection sensitivity even under interference. Its unobtrusiveness and model-agnostic nature make it practical for real-world applications, reducing detection and reverse engineering risks. Although our method can resist some model extraction attacks, it has limitations. Attackers might use jailbreak prompts to bypass preset system instructions, making it uncertain if our watermarking will be compromised. Defending against jailbreak prompts is complex and extensively studied, so we do not delve into whether system instructions will be executed. We hope ModelShield offers fresh insights for copyright protection of large language models. It provides a viable and efficient strategy for model owners to safeguard their IP rights at a low cost, fostering fairness and innovation in the LMaaS market.